### v1

# Educating LLMs like Human Students: Structure-aware Injection of Domain Knowledge

Abstract
This paper presents a pioneering methodology, termed StructTuning, to efficiently transform foundation Large Language Models (LLMs) into domain specialists. It significantly minimizes the training corpus requirement to a mere 0.3% while achieving an impressive 50% of traditional knowledge injection performance. Our method is inspired by the educational processes for human students, particularly how structured domain knowledge from textbooks is absorbed and then applied to tackle real-world challenges through specific exercises. Based on this, we propose a novel two-stage knowledge injection strategy: Structure-aware Continual Pre-Training (SCPT) and Structure-aware Supervised Fine-Tuning (SSFT). In the SCPT phase, we organize the training data into an auto-generated taxonomy of domain knowledge, enabling LLMs to effectively memorize textual segments linked to specific expertise within the taxonomy’s architecture. Subsequently, in the SSFT phase, we explicitly prompt models to reveal the underlying knowledge structure in their outputs, leveraging this structured domain insight to address practical problems adeptly. Our ultimate method has undergone extensive evaluations across model architectures and scales, using closed-book question-answering tasks on LongBench and MMedBench datasets. Remarkably, our method matches 50% of the improvement displayed by the state-of-the-art MMedLM2 on MMedBench, but with only 0.3% quantity of the training corpus. This breakthrough showcases the potential to scale up our StructTuning for stronger domain-specific LLMs. Code will be made public soon.
1 Introduction
Large language models (LLMs) have recently seen extensive deployment across various applications [42, 1, 20, 6]. When adapting foundational models (e.g., Llama series [39, 40, 2]) for specialized AI assistants in distinct domains [35, 13], developers usually employ two techniques to enhance LLMs’ proficiency: retrieval-augmented generation (RAG) [25] and domain knowledge injection [14]. While RAG effectively utilizes an external knowledge base to augment information, the retrieval process’s inherent noise poses challenges to generating reliable responses[55, 8]. Thus, another avenue of research focuses on injecting new knowledge into models via training techniques [11, 18, 30].
Continual pre-training [37, 19] has been preferred for integrating new, domain-specific knowledge into existing LLMs [9, 44, 35]. Nevertheless, it often involves costly auto-regressive training on billions of tokens aimlessly curated from the internet rather than from precise domain-specific textbooks [21]. For example, MMedLM [35] curates 25.5B tokens to derive a medical model, and DeepSeek-Coder [13] uses 2T tokens for coding adaptation. The common failure to learn effectively from limited textbook content has been attributed to insufficient data diversity [58], which however violates the observation during the human education process in Fig. 1: students gain knowledge by sequentially studying from textbooks, reviewing knowledge points and structures, and applying this knowledge through proper exercises. In this process, all the new data to learn is textbooks (structured content) and exercising examples (question-answering pairs), and students just adopt their world knowledge to memorize, understand, and apply the knowledge to become domain experts [23, 52].
Figure 1: Discrepancy between human learning process and vanilla LLM adaptation paradigm. Human students learn structured knowledge through textbooks section by section, with particular exercises on related knowledge points. Traditional LLM adaptation involves continual pre-training on data chunks from randomly concatenated text segments, with aimless supervised fine-tuning for conversation alignment. The inherent property of structured knowledge is ignored.
Inspired by this, we propose to inject the domain knowledge from textbooks into LLMs, like educating a human student, through a novel two-stage training strategy: Structure-aware Continual Pre-Training (SCPT) and Structure-aware Supervised Fine-Tuning (SSFT).
In the SCPT stage, we argue that high-quality textbook data alone can adequately infuse initial domain knowledge [12], where the organization of training corpora is crucial. Traditionally, text corpora are merely concatenated and divided into chunks of 2048 [35] or 4096 [13], while the inherent structure of the texts (e.g., catalogs of textbooks) is disregarded. Instead, we propose an automatic approach to maintain each chunk’s knowledge structure. We view each chunk as a knowledge point, employing advanced LLMs to extract domain knowledge taxonomy from the corpus efficiently, bypassing the need for manual annotation. Subsequently, LLMs are trained to predict textual content under the condition of linked knowledge points within the taxonomy, integrating individual training chunks with the entire knowledge framework. Ultimately, models are asked to memorize the knowledge structure to review the whole domain knowledge system.
In the SSFT stage, the goal shifts from knowledge injection to enabling LLMs to recall and utilize their acquired knowledge to tackle real-world challenges. We explicitly elicit knowledge structures in LLMs’ responses, as a beacon for models to targeted information retrieval or logical reasoning for reliable responses. To this end, we derive a scalable strategy to generate question-answer pairs as practice exercises by advanced LLMs such as GPT4 [1] or LLaMA3 [2]. In the scenarios with existing QA pairs like MMedBench [35], we retrieve related knowledge structures and content, instructing LLaMA3 to provide explanations from questions to answers based on these structures. For datasets lacking specific QA samples like LongBench [4], we randomly select knowledge branches from the domain taxonomy and prompt LLaMA3 to craft question-answer-explanation triplets for exercises.
Our ultimate approach, termed StructTuning, outperforms conventional methods in domain knowledge injection by emulating human learning processes through SCPT and SSFT phases. We extensively evaluate StructTuning’s effectiveness across different model architectures and sizes. For domain-adapted language models, we first examine their capability to recall injected knowledge through open-ended QA on the LongBench [4] dataset, then assess their application of this knowledge in addressing real-world issues via multiple-choice QA on the MMedBench dataset [35]. Results from both evaluations underscore the superiority of StructTuning. Remarkably, our strategy achieves a 50% improvement in knowledge injection compared to MMedLM2, the current benchmark in the medical domain, using merely 0.3% of the training data requirement. These findings underscore the potential to scale our method for enhancing domain-specific AI assistants significantly.
Our contribution is summarized as follows:
  • We proposed a novel two-stage training strategy, SCPT and SSFT, to inject domain knowledge into LLMs by preserving and utilizing the inherent structure of the training corpus.
  • We developed a scalable data construction framework to generate structure-aware training samples from original corpora, so as to facilitate the SCPT and SSFT stages.
  • We conducted extensive investigations on our StructTuning strategy on various data and model settings, and achieved competitive knowledge injection performance on MMedBench only using 0.3% of the training corpus.
2 Related Work
Domain Adaptation for Large Language Models. While pre-trained LLMs possess promising capabilities, their performance is often hampered by the scope and recency of their training data, which particularly affects smaller models in downstream applications [56, 43]. Continual Pre-Training (CPT) addresses this by perpetually updating a pre-trained model with domain-specific content [37, 50], with parameter-efficient tuning methods devised to curtail training costs [18, 29]. To keep pace with the latest information, models can be fine-tuned with supervised instruction-response pairs (SFT), thus staying current with the advancing knowledge landscape [30, 35]. Existing literature confirms that combining CPT and SFT is effective for LLMs to remain precise and up-to-date in dynamic fields like law [9, 32], finance [48, 26], medicine [44, 35], and coding [36, 13]. Our study builds upon this CPT-SFT framework, innovating with SCPT-SSFT strategies to efficiently and effectively infuse domain knowledge with the inherent structure hierarchy.
Conditional Language Modeling. The idea of continual pre-training language models on domain corpus in the condition of the knowledge structure is mainly inspired by CTRL [22]. Keskar et al. [22] demonstrates the effectiveness of steering text generation through control codes (one or two words) that signify the desired genre, style, or task. In the era of LLM, system prompt plays a similar role in controlling models’ responses to adapt to different needs and functionalities, such as role-playing, language style transfer, task setting, and behavior setting [7, 45, 3]. Our SCPT approach extends the control codes or system prompts to domain-specific knowledge structures, so as to guide the learning process and tailor the model’s output more closely to specialized fields.
Data Augmentation and Synthesis. Due to the lack of high-quality datasets, data augmentation has emerged as a promising solution to mimic real-world patterns [28]. Traditionally, these methods aim to artificially expand the training dataset size [49, 31] or generate entirely new samples that could help models learn better or adapt to specific tasks [38]. Yet, such methods often overlook the structured nature of domain knowledge, and the aimlessly generated samples may also lack diversity [33, 30], leading to potentially suboptimal training outcomes when applied for domain adaptations [30, 38]. Unlike these traditional approaches, our SSFT design is an innovative departure to address the challenge of retaining and utilizing the structured knowledge inherent in domain-specific content.
3 Methodology
This section presents our methodology to inject domain knowledge into pre-trained LLMs by utilizing the inherent knowledge structure, and Fig. 2 depicts our StructTuning framework. For a limited set of curated domain corpora (typically textbooks), we first develop an automatic approach to extract the knowledge structure, and associate the training chunks to the extracted knowledge structure and points (Sec. 3.1). Then, we design a two-stage training strategy to inject the highly structured domain knowledge into language models by mimicking the human education process, comprising the SCPT (Sec. 3.2) and SSFT (Sec. 3.3) techniques. The details are introduced as follows.
Figure 2: Framework for structure-aware knowledge injection. We extract the inherent knowledge structure in the training corpus, and associate training chunks to corresponding knowledge points. Models are continually pre-trained on data chunks in the condition of the knowledge structure, and fine-tuned with supervised QA samples to elicit their learned knowledge to solve real-world questions (including knowledge-intensive (KI) QA, 2- or multi-hop QA, etc.).
3.1 Automatic Extraction of Knowledge Structure
For web-crawled corpus, the meta-info of knowledge structures (e.g., the table content for a textbook) is usually inaccessible or inapplicable due to the data-parsing quality limitation, and all we have are those sequentially arranged text segments (e.g., page-by-page-chunked content). As shown in Fig. 2 (a), we aim to extract (or, recover) the knowledge structure from the raw corpus for subsequent domain knowledge injection, which is achieved by the following steps.
First, we use the spaCy library [^1] https://github.com/explosion/spaCy to split the content from a textbook at the sentence-level, and merge the sentences to form training chunks within a maximum size (e.g., 2048 tokens [35]). After that, we prompt the advanced Llama3-70B [2] model to summarize the title for each chunk, where the textual content with the abstractive title contributes to a “knowledge point”.
Then, we explore an automatic method to aggregate the sequential knowledge points and extract the inherent structure hierarchy by leveraging advanced language models. Particularly, we take the title list of those text chunks as the input (due to the limitation of context length) to instruct Llama3-70B [2] to identify the inherent structure within the title/chunk list, through in-context learning on the provided examples. The prompt template is displayed in Fig. 3, and the full version is in Appendix D.
Through the top-down structurization on the title list, we manage to extract the hierarchical knowledge structure from the original training corpus. Fig. 3 presents an example of the extracted structure, where we use the tree-like mindmap structure [46] to present the knowledge taxonomy from a textbook. The whole process does not involve human annotation, which reduces the cost and makes our method scalable for handling larger amounts of domain training corpora.
Figure 3: Left: prompt template to extract hierarchical knowledge structures from the given content. Right: example of representing the extracted knowledge structure by using a mindmap template.
After automatically extracting the domain knowledge structure and associating the original training chunks to related knowledge points, we delve into injecting domain knowledge through structure-aware continual pre-training (SCPT) and structure-aware supervised fine-tuning (SSFT).
3.2 Structure-aware Continual Pre-Training
In conventional knowledge injection methods, training corpora are randomly concatenated and chunked into text segments without distinguishing the original content, leading to the fact that models can only absorb domain knowledge that is emergent in the data diversity [33, 30, 35]. In this section, we present another solution to inject knowledge from limited pieces of textbooks by leveraging the highly abstractive and exhaustive domain knowledge structures for continual pre-training.
Figure 4: Example of prompt templates to bridge mindmap structure and textual contents.
We first transform the knowledge structure into natural languages using the same mindmap template in Fig. 3, and prepend it to each training chunk, forcing LLMs to memorize the textual content (knowledge points) in the condition of the associated knowledge path in the structure hierarchy. We collected 20 diversified templates from GPT-4 [1] to bridge mindmap structures and training chunks, one of which is displayed in Fig. 4. The prepended mindmap, as well as the template, does not produce auto-regressive loss. Losses are only calculated in the content part. Formally, we turn the original language modeling in vanilla CPT to conditioned modeling [22] in our SCPT stage:
$p({\bm{x}}^{k})=\prod_{i=1}^{n}{p(x^{k}_{i}|x^{k}_{<i})}\quad\Longrightarrow\quad p({\bm{x}}^{k}|{\bm{s}}^{k})=\prod_{i=1}^{n}{p(x^{k}_{i}|x^{k}_{<i},{\bm{s}}^{k})}$ (1)
where $p({\bm{x}}^{k})$ models the probability distribution for the $k$-th chunk ${\bm{x}}^{k}=(x^{k}_{1},\cdots,x^{k}_{n})$ via the chain rule of probability [5] on each token $x^{k}_{i}$, and ${\bm{s}}^{k}$ denotes the associated knowledge mindmap paths.
As illustrated in Fig. 2 (b), after traversing the $m$ knowledge points in extracted structures, models are asked to recall the whole knowledge hierarchy, i.e., to model the composed probability distribution:
$p(\bar{{\bm{s}}})=\prod_{k=1}^{m}{p({\bm{s}}^{k})}$ (2)
In SCPT, we mimic the human education process to inject knowledge into LLMs in a section-by-section manner, and replay the entire knowledge structure for the models to review and summarize the learned domain knowledge. These two steps iteratively alternate throughout training epochs.
Next, we will introduce how to teach LLMs to explicitly utilize their domain knowledge, which is learned in the SCPT stage, to solve practical problems by doing exercises with our SSFT technique.
3.3 Structure-aware Supervised Fine-Tuning
In traditional knowledge injection paradigms, supervised fine-tuning serves as a tool to align the (continually) pre-trained models to interactive chatbots, e.g., through massive question-answering exercises [9, 35]. However, most QA data augmentation or generation strategies aim at enlarging the quantity and enhancing the diversity of training samples [49, 31, 28], which neglects the nature of the highly structured domain knowledge. Therefore, our structure-aware supervised fine-tuning (SSFT) technique focuses on eliciting models’ structured knowledge learned during the SCPT stage, adapting LLMs to interactive and reliable domain experts.
Fig. 2 (c) illustrates the idea of synthesizing SSFT samples guided by domain knowledge structures, as extracted in Sec. 3.1. Specifically, we use the random walk algorithm to create knowledge paths with $1$ to $l$ branches in the original mindmap. For paths with only one knowledge point, we use the corresponding text content to prompt Llama3-70B [2] to generate knowledge-intensive question-answering pairs. For paths with two or more branches, we prompt Llama3-70B with the knowledge path and textual contents to synthesize 2- or multi-hop QA samples, which require specific reasoning along the knowledge structure to derive from questions to answers. Fig. 5 presents several examples.
Figure 5: QA samples synthesized for SSFT. We instruct Llama3-70B to generate (a) knowledge-intensive and (b) multi-hop questions and derive the diagnosis answers with explicit reasoning.
For every synthesized QA sample (${\bm{z}}$), we will prepend the relevant mindmap hierarchy to the answer, and add a CoT prompt in the question to construct another type of QA data (${\bm{z}}^{\prime}$) for SFT alignment. This design explicitly elicits the learned knowledge in models’ responses, teaching them how to apply the structured knowledge to address real-world problems. We use the two types of QA samples for training, as recommended by Qiu et al. [35]. During testing, we can either use the vanilla question as input to efficiently gather models’ answers, or take the CoT prompt to probe to what extent LLMs can memorize and leverage the injected knowledge to answer the questions.
Integrating with SCPT and SSFT, our StructTuning approach translates into remarkable efficacy and efficiency in domain knowledge injection, as comprehensively evaluated in the following sections.
4 Experiments
We design a comprehensive evaluation of our StructTuning through seveal experiments on two benchmarks. First, we investigate the free-form question-answering task on the LongBench [4] dataset, in order to verify the memorization and understanding of injected knowledge (the answer can be directly found in training corpora). Then, we delve into the multi-choice question-answering task on MMedBench [35], to explore how LLMs apply the injected knowledge in basic medicine to determine the real-world diagnosis for patients with logical reasoning.
4.1 Preliminary Investigation on Free-form Question-Answering
Datasets and Tasks. LongBench [4] is a multi-task benchmark tailored for open-book reading comprehension evaluation, where LLMs generate answers to given questions based on one or several lengthy passages. To focus on knowledge injection, we turn the open-book evaluation into a closed-book QA task, where contextual passages are injected into LLMs with training techniques, and models answer questions without input passages. We choose 7 subsets with 1,350 test examples from LongBench for single- and multi-document question-answering evaluation, and the remaining synthetic or code-orientated tasks are eliminated. More details are described in Sec. A.1.
Evaluation Metrics. For free-form QA in LongBench [4], we report the F1-Score between models’ outputs and ground-truth answers for both open-book and closed-book QA tasks. Besides, we also report the Recall for model responses to quantify the memorization degree of injected knowledge.
Investigated Models. For LongBench, we mainly investigate the knowledge injection to the Llama2-7B model [40], and report the Open- and Closed-Book QA performance for comparison.
Implementation Details. On LongBench for closed-book QA, we continually pre-train the Llama2-7B model on 10,476 passages for 3 epochs using a batch size of 128. The initial learning rate is 2e-5, decayed to 0 at the end of training with a cosine scheduler. Then, we query Llama3-70B to generate 2,700 SSFT QA examples, and eliminate the generated data whose answer shares over 0.5 F1-Score with one of the test samples to avoid data leakage. We train the knowledge-injected models for 1 epoch to avoid overfitting. Models are trained with the LLaMA-Factory framework  [^2] https://github.com/hiyouga/LLaMA-Factory.
Figure 6: An example for structure-aware responses.
Main Results. In Tab. 1, we first report the Open-Book QA (OBQA) baseline for Llama2-7B with passages inputs, which shows the best performance on MultiFieldQA (MFQA) [4] and 2WikiMultihopQA (2Wiki) [17] subsets. Then, we establish the Closed-Book QA (CBQA) baseline for pre-trained Llama2-7B, which aims to identify the knowledge that has already been equipped from the newly injected knowledge. Consequently, we follow the traditional CPT+SFT paradigm to inject passage contents into model parameters, and supplement the experiment of our SCPT+SSFT technique for comparison. According to the results shown in Tab. 1, the traditional continual pre-training strategy immediately improves the QA performance on six out of seven subsets (except the Qasper [10] subset), while the supervised fine-tuning alignment does not bring significant gains on the overall F1-Score measure. In contrast, our SCPT+SSFT approach successfully boosts the closed-book QA performance to 20.1% on average, even surpassing the open-book QA baseline of 18.7%. It indicates the vanilla SFT strategy can only regularize LLMs’ response styles, while our SSFT could teach LLMs to utilize their knowledge (injected in the SCPT stage) to answer corresponding questions, as exemplified in Fig. 6.
Table 1: F1 Score evaluation of Open-Book QA (OBQA) and Closed-Book QA (CBQA) tasks on the LongBench [4] dataset. The best results are marked in bold, and the secondary results are marked with underlines. The backbone model is Llama2-7B [40].
Task Adaptation SingleDoc-QA MultiDoc-QA Average
Qasper MFQA MFQAzh HpQA 2Wiki Musiq Duzh
OBQA - 19.2 36.8 11.9 25.4 32.8 9.4 5.2 18.7
CBQA - 17.3 15.9 1.4 5.7 6.7 3.3 1.8 7.4
CPT 16.8 23.6 17.5 18.0 18.6 6.7 14.3 16.5
CPT+SFT 16.8 23.1 13.2 21.3 19.1 10.4 13.4 16.8
SCPT 15.2 21.5 15.2 14.9 19.8 5.6 14.1 15.2
SCPT+SSFT 19.7 23.5 19.5 26.4 24.1 12.2 15.4 20.1
Table 2: Recall evaluation of the Closed-Book QA task. The best results are marked in bold.
Adaptation SingleDoc-QA MultiDoc-QA Average
Qasper MFQA MFQAzh HpQA 2Wiki Musiq Duzh
CPT 13.4 33.0 16.6 28.1 27.7 15.9 13.1 21.1
CPT+SFT 20.7 35.3 20.6 29.9 32.1 18.9 12.0 24.2
SCPT 18.8 42.5 17.7 35.7 36.4 20.5 15.3 26.7
SCPT+SSFT 30.5 44.6 24.3 40.8 42.0 21.8 16.8 31.5
To further support this claim, we compute Recall between model outputs and ground-truth answers in Tab. 2 as a proxy measure for memorizing and manipulating the injected knowledge [59]. Specifically, we use a CoT instruction of Answer the question and explain why. to elicit models’ knowledge in their responses. As shown in Tab. 2, our SCPT strategy has already achieved higher knowledge recall than the CPT-SFT paradigm (26.7% v.s. 24.2%). It implies our structure-aware continually pre-trained model has successfully associated the relevant passages with their entire knowledge structure for the given question, which provides the knowledge path to seek targeted information to derive the answer. As displayed in Fig. 6, our SSFT technique explicitly teaches the model to recall the learned knowledge and answer the questions, which further improves the knowledge recall in Tab. 2.
Figure 7: MindMap Recall
F1-Score BERTScore
0.61 0.87
In addition, we also statistically quantify the memorization of injected knowledge structures by comparing the mindmap in models’ responses with the ground-truth answers via lexical ROUGE-L [27] and semantic BERTScore [54] measures. The results shown in Fig. 7 indicate a relatively good memorization of the injected knowledge mindmap, showing the efficacy of our SCPT strategy.
4.2 In-depth Evaluation for Multi-choice QA Application
Datasets and Tasks. MMedBench [35] is a multilingual medical multi-choice QA benchmark, with 45,048 QA pairs for adapting LLMs to medical experts and 8,518 for testing. We collect 76M textbook corpora from MedTextBooks [21] and MMedC [35] for medical knowledge injection and use the training/test split from MMedBench for SFT/evaluation. Detailed setup is in Sec. A.2.
Evaluation Metrics. For the multi-choice QA task in MMedBench [35], we follow the default setting to calculate the accuracy on six language subsets, as well as the averaged scores.
Investigated Models. For MMedBench, we extend the investigated LLMs across model scales and architectures including Llama2-7B, Llama2-13B [40], and InternLM2-7B [57]. We also compare our knowledge-injected models with other open- or closed-source LLMs, such as the domain-specified MedAlpaca [15], ChatDoctor [53], PMC-LLaMA [47], and MMedLM [35] models.
Implementation Details. We first train LLMs for 3 epochs on medical textbooks with a batch size of 128. Then, we ask Llama3-70B to create structure-aware explanations for existing 45K QA samples in MMedBench’s training split and 33K extra entries by traversing extracted knowledge structures. Syntheses with overlapped options in the test set are removed. In the SFT phase, the learning rate is set as 1e-6 to avoid overfitting on such an amount of SFT samples, as suggested by Qiu et al. [35].
Main Results. In Tab. 3, we present the overall performance across a series of LLMs on the testing split of MMedBench. The results demonstrate the promising enhancement achieved by our StructTuning technique, which translates into consistently significant improvements on Llama2-7B (+8.78%), Llama2-13B (+6.17%), and InternLM2-7B (+4.46%). It shows the generalizability and scalability of our StructTuning strategy across model architectures and sizes. Notably, our method, comprising SCPT on 76M curated textbook corpora and SSFT on 78K QA samples, achieves 50% performance (4.46% v.s. 8.76%) against the state-of-the-art MMedLM2-7B model, which is trained on a huge MMedC [35] corpora of 25.5B tokens. Our approach shows much more efficiency in transforming pre-trained LLMs into domain experts by structured-aware knowledge injection, which largely reduces the cost for LLMs’ domain adaptation to 0.3%. However, our method is currently unable to achieve 100% of knowledge injection efficacy against MMedLM2, which may comprise two major reasons: (1) some of the knowledge in the training corpus has already been learned by the foundation InternLM2 model, and (2) the knowledge in our curated 76M training data cannot cover all the testing samples. We leave the promotion of achieving 100% improvement in our future work.
Table 3: Multiple-choice accuracy evaluation on MMedBench [35]. We report each model’s accuracy across six languages separately, with “Average” denoting the mean score over six languages. We also list out the data quantity for CPT and SFT stages during medical knowledge injection.
Model CPT SFT English Chinese Japanese French Russian Spanish Average
InternLM-7B - 45K 44.07 64.62 37.19 24.92 58.20 44.97 45.67 +0.00
MMedLM-7B 25B 45K 49.88 70.49 46.23 36.66 72.27 54.52 55.01 +9.34
InternLM2-7B - 45K 57.27 77.55 47.74 41.00 68.36 59.59 58.59 +0.00
MMedLM2-7B 25B 45K 61.74 80.01 61.81 52.09 80.47 67.65 67.30 +8.71
Llama2-7B - 45K 43.36 50.29 25.13 20.90 66.80 47.10 42.26 +0.00
+Ours 76M 78K 49.41 65.15 36.68 35.21 69.14 50.62 51.04 +8.78
Llama2-13B - 45K 51.37 57.97 32.66 25.08 69.92 52.99 48.33 +0.00
+Ours 76M 78K 53.02 68.30 37.78 41.71 70.70 55.51 54.50 +6.17
InternLM2-7B - 45K 57.27 77.55 47.74 41.00 68.36 59.59 58.59 +0.00
+Ours 76M 78K 60.80 79.19 50.75 45.34 75.39 66.85 63.05 +4.46
Ablation Studies. We conduct a comprehensive ablation study with the English split of the MMedBench dataset to evaluate our efficacy. Specifically, we take Llama2-7B as the backbone model, select the English textbooks [21] (with 26M tokens) for vanilla and our structure-aware continual pre-training, and use different QA samples for supervised fine-tuning. In particular, in Tab. 4, SFT refers to vanilla SFT with 10K English QA samples provided in MMedBench’s [35] training split, SSFT indicates structure-aware SFT on 10K training data, where the questions are the same as SFT while the answers are enhanced with knowledge structure explanation by Llama3-70B, as described in Sec. 3.3. SSFT* includes another 8K structure-aware QA syntheses by Llama3-70B, consisting of 18K entries for training. The training hyper-parameters follow the main experiment.
Results in Tab. 4 imply the vanilla CPT+SFT injection paradigm does bring considerable improvement of 1.73% in the English split, while our SCPT technique with vanilla SFT strategy presents a slightly higher accuracy of 46.50%. However, when combining SSFT with SCPT, our method immediately leads to a significant boost in the English split (49.96% v.s. 44.54%), demonstrating the efficacy and necessity of eliciting the learned structured knowledge to solve practical problems. Moreover, the supplementation of 8K extra QA syntheses surprisingly enhances the model performance on the other five subsets, resulting in a valuable gain on the average accuracy of 38.27%. This result demonstrates the knowledge transferability across different languages [24, 34]. After training with our SSFT strategy, LLMs can actively utilize the knowledge injected in one language to solve the problem in another language, further evidencing our superiority against the traditional SFT technique.
In addition, we observe that the commonly used RAG [25] strategy does not bring significant advantages to the MMedBench evaluation. The main reason lies in the gap between the pre-training corpus (comprising official knowledge statements from textbooks) and evaluated QA samples (originating from practical diagnosis records). To teach LLMs to apply relevant medical knowledge for reliable diagnosis answers, knowledge injection by CPT and SFT shows more advantages in this situation.
Table 4: Ablation studies with Llama2-7B on the English subset of MMedBench. The best results are marked in bold, and the secondary results are marked with underlines.
Adaptation English Chinese Japanese French Russian Spanish Average
- SFT 44.54 32.81 26.63 15.27 53.91 42.30 35.91
CPT SFT 46.27 32.57 26.13 17.36 50.00 40.63 35.49
SCPT SFT 46.50 32.14 20.10 18.17 53.91 39.97 35.13
SCPT SSFT 49.96 32.63 22.11 17.52 51.17 41.28 35.78
SCPT SSFT* 49.10 33.92 18.33 27.14 57.42 43.73 38.27
RAG 38.12 29.22 22.61 23.34 53.91 36.47 33.95
5 Conclusion
This work pioneers in incorporating structure-aware methodologies to enhance domain knowledge injection into large language models (LLMs). Through a novel two-stage training strategy, combining the SCPT and SSFT techniques, we have set a new precedent for efficiently adapting LLMs to specialized domains. The promising results and the scalable data construction framework we developed underscore the viability and potential of our method. The limitations and broader impacts are discussed in Appendix C. We hope to inspire further research in efficient and effective domain adaptation, moving a step closer to models that can truly emulate human intelligence.
Appendix A Implementation Details
A.1 Detailed Setup on LongBench
To focus on the investigation of knowledge injection, we choose 7 subsets from LongBench [4] across single- and multi-document QA tasks in English and Chinese, and the remaining synthetic or code-orientated tasks are eliminated:
  • Single-Doc QA. For single-document QA, we take three subsets from LongBench: (1) Qasper [10], featured by question-answering over NLP technical papers and annotated by NLP practitioners; (2) MultiFieldQA [4], manually curated from multiple data sources and annotated by Ph.D. students; and (3) MultiFieldQA-zh, the Chinese split also provided by Bai et al. [4], covering multiple Chinese scenarios. MultiFieldQA contains 150 Context-Question-Answer triplets to test, and the others adopted subsets include 200 pieces of test samples respectively.
  • Multi-Doc QA. Multi-document QA requires LLMs to extract and combine information from multiple documents to derive the answer, which is generally more challenging than single-doc QA. We take four multi-hop QA datasets: (1) HotpotQA [51], containing 2-hop questions written by native speakers given two related paragraphs; (2) 2WikiMultihopQA [17], involving up to 5-hop questions synthesized through manually designed templates on Wikipedia passages; (3) MuSiQue [41], carefully composed with up to 4-hop reasoning on an increased number of supporting and distracting context evidence; and (4) Dureader [16], developed based on Baidu Search and Baidu Zhidao and filtered by Bai et al. [4] to reduce the data noise. Each subset has 200 test samples.
A.2 Detailed Setup on MMedBench
Data for Evaluation. The Multilingual Medical Benchmark (MMedBench) [35] represents a comprehensive and diverse multilingual medical Question and Answering (QA) benchmark designed to evaluate models’ capabilities of understanding and processing medical content.
MMedBench’s robust dataset extends across 6 languages (i.e., English, Chinese, Japanese, French, Russian, and Spanish) and 21 medical fields, which include, but are not limited to, Internal Medicine, Biochemistry, Pharmacology, Psychiatry, and many others. It provides 45,048 training pairs and 8,518 testing pairs for diverse learning and testing scenarios. The training split is specifically designed for domain-specific finetuning of large language models (LLMs), while the entire testing set allows for a precise assessment of multi-choice question-answering performance. Statistics on six languages are displayed in Tab. A1. Notably, the benchmark includes scenarios where questions may have multiple correct answers (i.e., in Japanese and French subsets), introducing additional complexity for the model evaluation process.
Table A1: Sample statistics on MMedBench.
Split English Chinese Japanese French Russian Spanish Total
Train 10,178 27,400 1,590 2,171 1,052 2,657 45,048
Test 1,273 3,426 199 622 256 2,742 8,518
Data for Continual Pre-Training. To investigate high-quality domain knowledge injection for LLMs, we collect 18 English textbooks and 33 Chinese textbooks from the National Medical Board Examination in the USA and Mainland China, respectively [21]. All collected textbooks are originally in PDF format and Jin et al. [21] converted them into digital text via OCR and performed some clean-up pre-processing strategies to reduce the data noise. The English and Chinese textbooks count for around 26.1M and 21.5M tokens by Llama2 tokenizer [40]. Then, we randomly sample an extra 28M textbook corpora from MMedC [35] for the other languages (except the unavailable Japanese textbooks), and the final statistics are displayed in Tab. A2.
Table A2: Sample statistics on training data.
Stage English Chinese Japanese French Russian Spanish Total
CPT 26.1M 21.5M - 8.1M 10.3M 10.1M 76.1M
SFT 18.8K 39.1K 1.6K 5.3K 5.9K 7.5K 78.2K
Data for Supervised Fine-Tuning. As introduced in Sec. 3.3, we prompt Llama3-70B [2] to build the structure-aware answer explanations on top of the raw SFT samples in MMedBench’s training split, and generate extra QA pairs by traversing the extracted knowledge structure from textbooks. The final quantity statistics are presented in Tab. A2.
Appendix B Resource Requirement
We use 8 NVIDIA A100-80G GPUs to train all the models, and leverage 1-2 NVIDIA A100-80G GPUs for inference.
Appendix C Limitations and Discussions
Limitations. A noteworthy limitation of our work, despite its significant achievements, is centered around the exploration of the full potential of our StructTuning strategy for knowledge injection. While our method has successfully demonstrated a remarkable 50% improvement in domain-specific knowledge application and retention using merely 0.3% of the training corpus, the extent of data scalability needed to achieve 100% of the desired effectiveness remains unexplored. This limitation primarily arises from constraints in time and resources available during our research phase. We will delve into the investigations in our future work.
Broader Impacts. We discuss the positive and negative societal impacts as follows:
  • Positive Societal Impacts. Our StructTuning strategy could democratize access to advanced AI, enabling specialized AI assistants across various sectors such as healthcare and education. This method’s efficiency could improve diagnostic tools, personalized learning experiences, and enhanced decision-making processes. Furthermore, by mirroring the human learning process, our approach facilitates the development of AI systems that are easier for people to trust and interact with.
  • Negative Societal Impacts. The enhanced capabilities of LLMs might be exploited for generating convincing yet fraudulent content, raising concerns over misinformation and intellectual deceit. Additionally, if the structured knowledge base contains biases, it could inadvertently reinforce societal stereotypes, necessitating rigorous curation to ensure fairness and accuracy. To counteract potential misuse, controlled model releases and continuous monitoring for abuse are recommended. Ensuring the diversity and contemporaneity of training data can mitigate bias. Collaborative efforts with domain experts, ethicists, and policymakers are vital in developing guidelines for the technology’s ethical use.
Appendix D Prompt Template for Structurization
Fig. A1 displays the few-shot prompt template to query giant commercial LLMs to execute structurization. We first introduce the instruction for structurization (including the task definition and output formats), then provide two representative examples to further illustrate the process. The first example in Fig. A2 tells the model to focus on the existing numerical or enumeration indicators to assist in constructing the aspect-description structure. The second example in Fig. A3 teaches the model to automatically summarize the aspects and attach their corresponding descriptions from the raw text sequences. Finally, in Fig. A1 we ask commercial LLMs to structurize the $input\_statement$ from user input as expected.
Figure A1: Prompt template for few-shot structurization.
Figure A2: The first example for few-shot structurization.
Figure A3: The second example for few-shot structurization.

### v2

@@
-Educating LLMs like Human Students: Structure-aware Injection of Domain Knowledge
+Structure-aware Domain Knowledge Injection for Large Language Models
 Abstract
-This paper presents a pioneering methodology, termed StructTuning, to efficiently transform foundation Large Language Models (LLMs) into domain specialists.
-It significantly minimizes the training corpus requirement to a mere 0.3% while achieving an impressive 50% of traditional knowledge injection performance.
-Our method is inspired by the educational processes for human students, particularly how structured domain knowledge from textbooks is absorbed and then applied to tackle real-world challenges through specific exercises.
-Based on this, we propose a novel two-stage knowledge injection strategy: Structure-aware Continual Pre-Training (SCPT) and Structure-aware Supervised Fine-Tuning (SSFT).
-In the SCPT phase, we organize the training data into an auto-generated taxonomy of domain knowledge, enabling LLMs to effectively memorize textual segments linked to specific expertise within the taxonomy’s architecture.
-Subsequently, in the SSFT phase, we explicitly prompt models to reveal the underlying knowledge structure in their outputs, leveraging this structured domain insight to address practical problems adeptly.
+This paper introduces a pioneering methodology, termed StructTuning, to efficiently transform foundation Large Language Models (LLMs) into domain specialists.
+It significantly reduces the training corpus requirement to a mere 0.3%, while achieving an impressive 50% of traditional knowledge injection performance.
+Our method is inspired by the educational processes of human students, particularly how structured domain knowledge from textbooks is assimilated and subsequently applied to tackle real-world challenges through specific exercises.
+Based on this, we propose a novel two-stage strategy for knowledge injection and alignment: Structure-aware Continual Pre-Training (SCPT) and Structure-aware Supervised Fine-Tuning (SSFT).
+In the SCPT phase, we automatically extract the domain knowledge taxonomy and reorganize the training corpora, enabling LLMs to effectively link textual segments to targeted knowledge points within the taxonomy.
+In the SSFT phase, we explicitly prompt models to elucidate the underlying knowledge structure in their outputs, leveraging the structured domain insight to address practical problems.
 Our ultimate method has undergone extensive evaluations across model architectures and scales, using closed-book question-answering tasks on LongBench and MMedBench datasets.
-Remarkably, our method matches 50% of the improvement displayed by the state-of-the-art MMedLM2 on MMedBench, but with only 0.3% quantity of the training corpus.
-This breakthrough showcases the potential to scale up our StructTuning for stronger domain-specific LLMs.
-Code will be made public soon.
+Remarkably, our method demonstrates the potential of comparable improvement against the state-of-the-art MMedLM2 on MMedBench, while significantly reducing the training costs to 5%.
+This breakthrough paves the way for scaling up our StructTuning for stronger domain-specific LLMs with comprehensive data utilization.
+Code is available at https://github.com/alibaba/struxgpt.
 1 Introduction
-Large language models (LLMs) have recently seen extensive deployment across various applications [42, 1, 20, 6].
-When adapting foundational models (e.g., Llama series [39, 40, 2]) for specialized AI assistants in distinct domains [35, 13], developers usually employ two techniques to enhance LLMs’ proficiency: retrieval-augmented generation (RAG) [25] and domain knowledge injection [14].
-While RAG effectively utilizes an external knowledge base to augment information, the retrieval process’s inherent noise poses challenges to generating reliable responses[55, 8].
-Thus, another avenue of research focuses on injecting new knowledge into models via training techniques [11, 18, 30].
-Continual pre-training [37, 19] has been preferred for integrating new, domain-specific knowledge into existing LLMs [9, 44, 35].
-Nevertheless, it often involves costly auto-regressive training on billions of tokens aimlessly curated from the internet rather than from precise domain-specific textbooks [21].
-For example, MMedLM [35] curates 25.5B tokens to derive a medical model, and DeepSeek-Coder [13] uses 2T tokens for coding adaptation.
-The common failure to learn effectively from limited textbook content has been attributed to insufficient data diversity [58], which however violates the observation during the human education process in Fig.
+Large language models (LLMs) have recently seen extensive deployment across various applications (Vaswani et al., 2017; Achiam et al., 2023; Jiang et al., 2023; Bi et al., 2024).
+When adapting foundational models (e.g., Llama series (Touvron et al., 2023a; b; Dubey et al., 2024)) for specialized AI assistants in distinct domains (Qiu et al., 2024; Guo et al., 2024), developers usually employ two techniques to enhance LLMs’ proficiency: retrieval-augmented generation (RAG) (Lewis et al., 2020) and domain knowledge injection (Gururangan et al., 2020).
+While RAG effectively utilizes an external knowledge base to augment information, the retrieval process’s inherent noise poses challenges to generating reliable responses, especially in scenarios requiring logical reasoning where there is a semantic gap between the user’s query and the knowledge base.(Zhang et al., 2023; Chen et al., 2023).
+Thus, another avenue of research focuses on injecting new knowledge into models via training techniques (Gu et al., 2021; Hu et al., 2021; Mecklenburg et al., 2024).
+Continual pre-training (Sun et al., 2020; Ibrahim et al., 2024) has been preferred for integrating new, domain-specific knowledge into existing LLMs (Cui et al., 2023; Wang et al., 2023b; Qiu et al., 2024).
+Nevertheless, it often entails resource-intensive auto-regressive training on billions of tokens from the internet to learn fragmented knowledge points, rather than absorbing structured knowledge from a few domain-specific textbooks (Jin et al., 2020).
+For example, MMedLM (Qiu et al., 2024) curates 25.5B tokens to derive a medical model, and DeepSeek-Coder (Guo et al., 2024) uses 2T tokens for coding adaptation.
+The common failure to learn effectively from limited textbook content has been attributed to insufficient data diversity (Zhu & Li, 2023a), which however violates the observation during the human education process in Fig.
 1: students gain knowledge by sequentially studying from textbooks, reviewing knowledge points and structures, and applying this knowledge through proper exercises.
-In this process, all the new data to learn is textbooks (structured content) and exercising examples (question-answering pairs), and students just adopt their world knowledge to memorize, understand, and apply the knowledge to become domain experts [23, 52].
+In this process, all the new data to learn are textbooks (structured content) and exercising examples (question-answering pairs), and students just adopt their world knowledge to memorize, understand, and apply the knowledge to become domain experts (Krathwohl, 2002; Yu et al., 2023).
 Figure 1: Discrepancy between human learning process and vanilla LLM adaptation paradigm.
@@
 The inherent property of structured knowledge is ignored.
-Inspired by this, we propose to inject the domain knowledge from textbooks into LLMs, like educating a human student, through a novel two-stage training strategy: Structure-aware Continual Pre-Training (SCPT) and Structure-aware Supervised Fine-Tuning (SSFT).
-In the SCPT stage, we argue that high-quality textbook data alone can adequately infuse initial domain knowledge [12], where the organization of training corpora is crucial.
-Traditionally, text corpora are merely concatenated and divided into chunks of 2048 [35] or 4096 [13], while the inherent structure of the texts (e.g., catalogs of textbooks) is disregarded.
+Inspired by this, we propose to inject the domain knowledge from textbooks into LLMs, as educating a human student, through a novel two-stage training strategy: Structure-aware Continual Pre-Training (SCPT) and Structure-aware Supervised Fine-Tuning (SSFT).
+In the SCPT stage, we argue that high-quality textbook data can adequately infuse the domain knowledge (Gunasekar et al., 2023), where the organization of training corpora is crucial.
+In conventional paradigms, as illustrated in Fig.
+1, text corpora are simply concatenated and divided into chunks of 2048 (Qiu et al., 2024) or 4096 (Guo et al., 2024), while the inherent structure of the texts (e.g., catalogs of textbooks) is disregarded.
 Instead, we propose an automatic approach to maintain each chunk’s knowledge structure.
-We view each chunk as a knowledge point, employing advanced LLMs to extract domain knowledge taxonomy from the corpus efficiently, bypassing the need for manual annotation.
-Subsequently, LLMs are trained to predict textual content under the condition of linked knowledge points within the taxonomy, integrating individual training chunks with the entire knowledge framework.
-Ultimately, models are asked to memorize the knowledge structure to review the whole domain knowledge system.
+We view each chunk as a knowledge point, and employ advanced LLMs to efficiently extract domain knowledge taxonomy from the corpus, bypassing the need for manual annotation.
+Subsequently, LLMs are trained to predict the textual content (corresponding to the knowledge point) under the condition of the knowledge path within the domain structure, linking individual training chunks with the entire knowledge architecture.
+Finally, models are asked to memorize the knowledge structure to review the whole domain knowledge system.
 In the SSFT stage, the goal shifts from knowledge injection to enabling LLMs to recall and utilize their acquired knowledge to tackle real-world challenges.
-We explicitly elicit knowledge structures in LLMs’ responses, as a beacon for models to targeted information retrieval or logical reasoning for reliable responses.
-To this end, we derive a scalable strategy to generate question-answer pairs as practice exercises by advanced LLMs such as GPT4 [1] or LLaMA3 [2].
-In the scenarios with existing QA pairs like MMedBench [35], we retrieve related knowledge structures and content, instructing LLaMA3 to provide explanations from questions to answers based on these structures.
-For datasets lacking specific QA samples like LongBench [4], we randomly select knowledge branches from the domain taxonomy and prompt LLaMA3 to craft question-answer-explanation triplets for exercises.
+We explicitly elicit knowledge paths in LLMs’ responses, as a beacon for models to targeted information retrieval or logical reasoning for reliable responses.
+To this end, we derive a scalable strategy to generate question-answer pairs as practice exercises by powerful LLMs such as GPT4 (Achiam et al., 2023) or LLaMA3 (Dubey et al., 2024).
+In the scenarios with existing QA pairs like MMedBench (Qiu et al., 2024), we retrieve related knowledge structure and content, instructing LLaMA3 to provide explanations from questions to answers based on the knowledge paths.
+For datasets lacking specific QA samples like LongBench (Bai et al., 2023), we randomly select knowledge paths from the domain taxonomy and prompt LLaMA3 to craft question-answer-explanation triplets for training exercises.
 Our ultimate approach, termed StructTuning, outperforms conventional methods in domain knowledge injection by emulating human learning processes through SCPT and SSFT phases.
 We extensively evaluate StructTuning’s effectiveness across different model architectures and sizes.
-For domain-adapted language models, we first examine their capability to recall injected knowledge through open-ended QA on the LongBench [4] dataset, then assess their application of this knowledge in addressing real-world issues via multiple-choice QA on the MMedBench dataset [35].
-Results from both evaluations underscore the superiority of StructTuning.
-Remarkably, our strategy achieves a 50% improvement in knowledge injection compared to MMedLM2, the current benchmark in the medical domain, using merely 0.3% of the training data requirement.
-These findings underscore the potential to scale our method for enhancing domain-specific AI assistants significantly.
+For domain-adapted language models, we first examine their capability to recall the injected knowledge through open-ended QA on the LongBench (Bai et al., 2023) dataset, then assess their application of this knowledge in addressing real-world issues via multiple-choice QA on MMedBench (Qiu et al., 2024).
+Both evaluations underscore the superiority of StructTuning.
+Remarkably, we achieve a 50% improvement in knowledge injection compared to the SOTA MMedLM2 in the medical domain, using merely 0.3% of the training data requirement.
+And StructTuning illustrates the potential of comparable performance with only 5% of training costs.
+These findings reveal the scalability of our method for enhancing domain-specific AI assistants with further comprehensive data utilization.
 Our contribution is summarized as follows:
@@
   • We developed a scalable data construction framework to generate structure-aware training samples from original corpora, so as to facilitate the SCPT and SSFT stages.
-  • We conducted extensive investigations on our StructTuning strategy on various data and model settings, and achieved competitive knowledge injection performance on MMedBench only using 0.3% of the training corpus.
+  • We conducted extensive investigations on our StructTuning strategy on various data and model settings, and comprehensively illustrate our superiority in knowledge injection.
 2 Related Work
 Domain Adaptation for Large Language Models.
-While pre-trained LLMs possess promising capabilities, their performance is often hampered by the scope and recency of their training data, which particularly affects smaller models in downstream applications [56, 43].
-Continual Pre-Training (CPT) addresses this by perpetually updating a pre-trained model with domain-specific content [37, 50], with parameter-efficient tuning methods devised to curtail training costs [18, 29].
-To keep pace with the latest information, models can be fine-tuned with supervised instruction-response pairs (SFT), thus staying current with the advancing knowledge landscape [30, 35].
-Existing literature confirms that combining CPT and SFT is effective for LLMs to remain precise and up-to-date in dynamic fields like law [9, 32], finance [48, 26], medicine [44, 35], and coding [36, 13].
+While pre-trained LLMs possess promising capabilities, their performance is often hampered by the scope and recency of their training data, which particularly affects smaller models in downstream applications (Zhao et al., 2023; Wang et al., 2023a).
+Continual Pre-Training (CPT) addresses this by perpetually updating a pre-trained model with domain-specific content (Sun et al., 2020; Xu et al., 2023b).
+To keep pace with the latest information, models can be fine-tuned with supervised instruction-response pairs (SFT), thus staying current with the advancing knowledge landscape (Mecklenburg et al., 2024; Qiu et al., 2024).
+Existing literature confirms that combining CPT and SFT is effective for LLMs to remain precise and up-to-date in dynamic fields like law (Cui et al., 2023; Nguyen, 2023) and medicine (Wang et al., 2023b; Qiu et al., 2024).
 Our study builds upon this CPT-SFT framework, innovating with SCPT-SSFT strategies to efficiently and effectively infuse domain knowledge with the inherent structure hierarchy.
-Conditional Language Modeling.
-The idea of continual pre-training language models on domain corpus in the condition of the knowledge structure is mainly inspired by CTRL [22].
-Keskar et al.
-[22] demonstrates the effectiveness of steering text generation through control codes (one or two words) that signify the desired genre, style, or task.
-In the era of LLM, system prompt plays a similar role in controlling models’ responses to adapt to different needs and functionalities, such as role-playing, language style transfer, task setting, and behavior setting [7, 45, 3].
-Our SCPT approach extends the control codes or system prompts to domain-specific knowledge structures, so as to guide the learning process and tailor the model’s output more closely to specialized fields.
+Structure-aware Knowledge Aggregation.
+A branch of researchers follows the conventional paradigm to extract entity-relation-entity triplets from texts to construct knowledge graphs (Pan et al., 2024), to enhance LLMs’s factual knowledge and logical reasoning by feature aggregation (Liu et al., 2020; Zhang et al., 2022), prompt engineering (Wen et al., 2023; Wang et al., 2023c), training data synthesis (Tang et al., 2024), etc.
+In these cases, each node corresponds to either a specific entity or an abstract concept, lacking the capability to present an informative and self-contained knowledge point.
+Some works have recently related a piece of descriptive text to a knowledge point, and constructed the knowledge structure for LLMs’ retrieval-augmented generation (Sarthi et al., 2024; Dong et al., 2024), where the top-to-down retrieval provides precise information-seeking paths along the knowledge structure.
+In this paper, we extend the structure-aware knowledge aggregation to LLMs’ training phase, injecting the whole domain knowledge structure into LLMs’ by linking the training samples into corresponding knowledge points and reasoning paths.
 Data Augmentation and Synthesis.
-Due to the lack of high-quality datasets, data augmentation has emerged as a promising solution to mimic real-world patterns [28].
-Traditionally, these methods aim to artificially expand the training dataset size [49, 31] or generate entirely new samples that could help models learn better or adapt to specific tasks [38].
-Yet, such methods often overlook the structured nature of domain knowledge, and the aimlessly generated samples may also lack diversity [33, 30], leading to potentially suboptimal training outcomes when applied for domain adaptations [30, 38].
-Unlike these traditional approaches, our SSFT design is an innovative departure to address the challenge of retaining and utilizing the structured knowledge inherent in domain-specific content.
+Traditional methods aim to artificially expand the training dataset size (Xu et al., 2023a; Mukherjee et al., 2023) or generate entirely new samples that could help models learn better or adapt to specific tasks (Tang et al., 2024).
+Yet, they often overlook the structured nature of domain knowledge, and the aimlessly generated samples may also lack diversity (Ovadia et al., 2023; Mecklenburg et al., 2024), leading to potentially suboptimal training outcomes when applied for domain adaptations (Mecklenburg et al., 2024; Tang et al., 2024).
+By contrast, our SSFT design is an innovative departure to address the challenge of retaining and utilizing the structured knowledge inherent in domain-specific content.
 3 Methodology
-This section presents our methodology to inject domain knowledge into pre-trained LLMs by utilizing the inherent knowledge structure, and Fig.
-2 depicts our StructTuning framework.
-For a limited set of curated domain corpora (typically textbooks), we first develop an automatic approach to extract the knowledge structure, and associate the training chunks to the extracted knowledge structure and points (Sec.
+Fig.
+2 depicts our StructTuning methodology to inject domain knowledge into pre-trained LLMs using the inherent knowledge structure.
+With curated domain corpora (typically a few textbooks), we first develop an automatic approach to extract the knowledge structure, and associate text chunks to corresponding knowledge paths and points (Sec.
 3.1).
@@
 3.3) techniques.
-The details are introduced as follows.
 Figure 2: Framework for structure-aware knowledge injection.
@@
 3.1 Automatic Extraction of Knowledge Structure
-For web-crawled corpus, the meta-info of knowledge structures (e.g., the table content for a textbook) is usually inaccessible or inapplicable due to the data-parsing quality limitation, and all we have are those sequentially arranged text segments (e.g., page-by-page-chunked content).
+For web-crawled corpus, previous data pre-processing focuses on quality assessment for individual documents (Bi et al., 2024), while the meta-info of knowledge structures (e.g., the table content for a textbook) is usually neglected or filtered out, and all we have are those sequentially arranged text segments (e.g., page-by-page-chunked content).
 As shown in Fig.
-2 (a), we aim to extract (or, recover) the knowledge structure from the raw corpus for subsequent domain knowledge injection, which is achieved by the following steps.
-First, we use the spaCy library [^1] https://github.com/explosion/spaCy to split the content from a textbook at the sentence-level, and merge the sentences to form training chunks within a maximum size (e.g., 2048 tokens [35]).
-After that, we prompt the advanced Llama3-70B [2] model to summarize the title for each chunk, where the textual content with the abstractive title contributes to a “knowledge point”.
-Then, we explore an automatic method to aggregate the sequential knowledge points and extract the inherent structure hierarchy by leveraging advanced language models.
-Particularly, we take the title list of those text chunks as the input (due to the limitation of context length) to instruct Llama3-70B [2] to identify the inherent structure within the title/chunk list, through in-context learning on the provided examples.
-The prompt template is displayed in Fig.
-3, and the full version is in Appendix D.
-Through the top-down structurization on the title list, we manage to extract the hierarchical knowledge structure from the original training corpus.
+2 (a), we aim to extract (or, recover) the knowledge structure from the raw corpus for subsequent domain knowledge injection.
+First, we use spaCy [^1] https://github.com/explosion/spaCy to split the content from a textbook at the paragraph-level, and merge the sentences to form training chunks within a maximum size (e.g., 2048 tokens (Qiu et al., 2024)).
+After that, we prompt the advanced Llama3-70B (Dubey et al., 2024) model to summarize the title for each chunk, where the textual content with the abstractive title jointly contributes to a “knowledge point”.
+Then, we automatically aggregate knowledge points and extract the inherent structure hierarchy by leveraging advanced language models.
+Inspired by Liu et al.
+(2024a), we take the title list to instruct a specifically developed 7B model to identify the inherent knowledge structure within the text chunks.
+A prompt example is displayed in Fig.
+3, and the detailed implementation is presented in Sec.
+B.1.
+In particular, Sec.
+B.1 and Sec.
+B.2 verify that our specialized 7B model can identify sufficiently precise knowledge structure for effective and efficient domain adaptation, as more powerful LLMs like LLaMA3-70B (Dubey et al., 2024) and GPT-3.5 (Brown et al., 2020) cannot bring significant enhancement while largely increase the inference costs.
 Fig.
-3 presents an example of the extracted structure, where we use the tree-like mindmap structure [46] to present the knowledge taxonomy from a textbook.
-The whole process does not involve human annotation, which reduces the cost and makes our method scalable for handling larger amounts of domain training corpora.
-Figure 3: Left: prompt template to extract hierarchical knowledge structures from the given content. Right: example of representing the extracted knowledge structure by using a mindmap template.
+3 presents an example of the extracted structure, where we use the tree-like mindmap structure (Wen et al., 2023) to present the knowledge taxonomy from a textbook.
+And Fig.
+A1 showcases how to deal with non-textbook data.
+The whole process does not involve human annotation, which reduces the cost and makes our method scalable for larger domain training corpora.
+Figure 3: Left: prompt template to extract hierarchical knowledge structures from the given content. Right: example of the extracted knowledge structure presented by a mindmap format.
 After automatically extracting the domain knowledge structure and associating the original training chunks to related knowledge points, we delve into injecting domain knowledge through structure-aware continual pre-training (SCPT) and structure-aware supervised fine-tuning (SSFT).
 3.2 Structure-aware Continual Pre-Training
-In conventional knowledge injection methods, training corpora are randomly concatenated and chunked into text segments without distinguishing the original content, leading to the fact that models can only absorb domain knowledge that is emergent in the data diversity [33, 30, 35].
+In conventional knowledge injection methods, training corpora are randomly concatenated and chunked into text segments without distinguishing the original content, leading to the fact that models can only absorb domain knowledge that is emergent in the data diversity (Ovadia et al., 2023; Mecklenburg et al., 2024; Qiu et al., 2024).
 In this section, we present another solution to inject knowledge from limited pieces of textbooks by leveraging the highly abstractive and exhaustive domain knowledge structures for continual pre-training.
@@
 3, and prepend it to each training chunk, forcing LLMs to memorize the textual content (knowledge points) in the condition of the associated knowledge path in the structure hierarchy.
-We collected 20 diversified templates from GPT-4 [1] to bridge mindmap structures and training chunks, one of which is displayed in Fig.
+We collected 20 diversified templates from GPT-4 (Achiam et al., 2023) to bridge mindmap structures and training chunks, one of which is displayed in Fig.
 4.
@@
 Losses are only calculated in the content part.
-Formally, we turn the original language modeling in vanilla CPT to conditioned modeling [22] in our SCPT stage:
+Formally, we turn the original language modeling in vanilla CPT to conditioned modeling (Keskar et al., 2019) in our SCPT stage:
 $p({\bm{x}}^{k})=\prod_{i=1}^{n}{p(x^{k}_{i}|x^{k}_{<i})}\quad\Longrightarrow\quad p({\bm{x}}^{k}|{\bm{s}}^{k})=\prod_{i=1}^{n}{p(x^{k}_{i}|x^{k}_{<i},{\bm{s}}^{k})}$ (1)
-where $p({\bm{x}}^{k})$ models the probability distribution for the $k$-th chunk ${\bm{x}}^{k}=(x^{k}_{1},\cdots,x^{k}_{n})$ via the chain rule of probability [5] on each token $x^{k}_{i}$, and ${\bm{s}}^{k}$ denotes the associated knowledge mindmap paths.
+where $p({\bm{x}}^{k})$ models the probability distribution for the $k$-th chunk ${\bm{x}}^{k}=(x^{k}_{1},\cdots,x^{k}_{n})$ via the chain rule of probability (Bengio et al., 2000) on each token $x^{k}_{i}$, and ${\bm{s}}^{k}$ denotes the associated knowledge mindmap paths.
 As illustrated in Fig.
@@
 3.3 Structure-aware Supervised Fine-Tuning
-In traditional knowledge injection paradigms, supervised fine-tuning serves as a tool to align the (continually) pre-trained models to interactive chatbots, e.g., through massive question-answering exercises [9, 35].
-However, most QA data augmentation or generation strategies aim at enlarging the quantity and enhancing the diversity of training samples [49, 31, 28], which neglects the nature of the highly structured domain knowledge.
+In traditional knowledge injection paradigms, supervised fine-tuning serves as a tool to align the (continually) pre-trained models to interactive chatbots, e.g., through massive question-answering exercises (Cui et al., 2023; Qiu et al., 2024).
+However, most QA data augmentation or generation strategies aim at enlarging the quantity and enhancing the diversity of training samples (Xu et al., 2023a; Mukherjee et al., 2023; Liu et al., 2024b), which neglects the nature of the highly structured domain knowledge.
 Therefore, our structure-aware supervised fine-tuning (SSFT) technique focuses on eliciting models’ structured knowledge learned during the SCPT stage, adapting LLMs to interactive and reliable domain experts.
 Fig.
-2 (c) illustrates the idea of synthesizing SSFT samples guided by domain knowledge structures, as extracted in Sec.
-3.1.
+2 (c) illustrates the idea of synthesizing SSFT samples guided by domain knowledge structures.
 Specifically, we use the random walk algorithm to create knowledge paths with $1$ to $l$ branches in the original mindmap.
-For paths with only one knowledge point, we use the corresponding text content to prompt Llama3-70B [2] to generate knowledge-intensive question-answering pairs.
+For paths linking to a single knowledge point, we use the corresponding text content to prompt Llama3-70B (Dubey et al., 2024) to generate knowledge-intensive question-answering pairs.
 For paths with two or more branches, we prompt Llama3-70B with the knowledge path and textual contents to synthesize 2- or multi-hop QA samples, which require specific reasoning along the knowledge structure to derive from questions to answers.
@@
 We use the two types of QA samples for training, as recommended by Qiu et al.
-[35].
+(2024).
 During testing, we can either use the vanilla question as input to efficiently gather models’ answers, or take the CoT prompt to probe to what extent LLMs can memorize and leverage the injected knowledge to answer the questions.
@@
 4 Experiments
-We design a comprehensive evaluation of our StructTuning through seveal experiments on two benchmarks.
-First, we investigate the free-form question-answering task on the LongBench [4] dataset, in order to verify the memorization and understanding of injected knowledge (the answer can be directly found in training corpora).
-Then, we delve into the multi-choice question-answering task on MMedBench [35], to explore how LLMs apply the injected knowledge in basic medicine to determine the real-world diagnosis for patients with logical reasoning.
+We design a comprehensive evaluation of our StructTuning through several experiments on two benchmarks.
+First, we investigate the free-form question-answering task on the LongBench (Bai et al., 2023) dataset, in order to verify the memorization and understanding of injected knowledge (the answer can be directly found in training corpora).
+Then, we delve into the multi-choice question-answering task on MMedBench (Qiu et al., 2024), to explore how LLMs apply the injected knowledge in basic medicine to determine the real-world diagnosis for patients with logical reasoning.
 4.1 Preliminary Investigation on Free-form Question-Answering
+Table 1: Recall evaluation of Open-Book QA (OBQA) and Closed-Book QA (CBQA) tasks on the LongBench (Bai et al., 2023) dataset.
+The best results are marked in bold, and the secondary results are marked with underlines.
+The backbone model is Llama2-7B (Touvron et al., 2023b).
+Task Adaptation SingleDoc-QA MultiDoc-QA Average
+Qasper MFQA MFQAzh HpQA 2Wiki Musiq Duzh
+OBQA - 39.7 44.3 17.5 30.1 35.5 11.9 9.6 26.9
+CBQA CPT+SFT 20.7 35.3 20.6 29.9 32.1 18.9 12.0 24.2
+SCPT+SFT 18.8 42.5 17.7 35.7 36.4 20.5 15.3 26.7
+SCPT+SSFT 30.5 44.6 24.3 40.8 42.0 21.8 16.8 31.5
 Datasets and Tasks.
-LongBench [4] is a multi-task benchmark tailored for open-book reading comprehension evaluation, where LLMs generate answers to given questions based on one or several lengthy passages.
-To focus on knowledge injection, we turn the open-book evaluation into a closed-book QA task, where contextual passages are injected into LLMs with training techniques, and models answer questions without input passages.
+LongBench (Bai et al., 2023) is a multi-task benchmark tailored for open-book reading comprehension evaluation, where LLMs generate answers to given questions based on one or several input passages.
+To focus on knowledge injection, we turn the open-book evaluation into a closed-book QA task, where LLMs are trained on contextual passages and queried with questions only.
 We choose 7 subsets with 1,350 test examples from LongBench for single- and multi-document question-answering evaluation, and the remaining synthetic or code-orientated tasks are eliminated.
@@
 Evaluation Metrics.
-For free-form QA in LongBench [4], we report the F1-Score between models’ outputs and ground-truth answers for both open-book and closed-book QA tasks.
-Besides, we also report the Recall for model responses to quantify the memorization degree of injected knowledge.
-Investigated Models. For LongBench, we mainly investigate the knowledge injection to the Llama2-7B model [40], and report the Open- and Closed-Book QA performance for comparison.
+To quantify the knowledge memorization degree, we mainly report the recall (Zhu & Li, 2023b) for models’ outputs against ground-truth answers for the free-form QA tasks.
+In Sec.
+B.3 we also evaluate our method with the F1-score measure for a thorough comparison.
+Investigated Models. For LongBench, we mainly investigate the knowledge injection to Llama2-7B (Touvron et al., 2023b) to compare the open- and closed-book QA performance.
 Implementation Details.
-On LongBench for closed-book QA, we continually pre-train the Llama2-7B model on 10,476 passages for 3 epochs using a batch size of 128.
+For closed-book QA, the Llama2-7B model is continually pre-trained on 10,476 passages for 3 epochs using a batch size of 128.
 The initial learning rate is 2e-5, decayed to 0 at the end of training with a cosine scheduler.
-Then, we query Llama3-70B to generate 2,700 SSFT QA examples, and eliminate the generated data whose answer shares over 0.5 F1-Score with one of the test samples to avoid data leakage.
+Then, we query Llama3-70B to generate 2,700 SSFT QA examples and remove those with over 0.5 F1-Score similarity to test samples to prevent data leakage.
 We train the knowledge-injected models for 1 epoch to avoid overfitting.
-Models are trained with the LLaMA-Factory framework  [^2] https://github.com/hiyouga/LLaMA-Factory.
 Figure 6: An example for structure-aware responses.
 Main Results.
-In Tab.
-1, we first report the Open-Book QA (OBQA) baseline for Llama2-7B with passages inputs, which shows the best performance on MultiFieldQA (MFQA) [4] and 2WikiMultihopQA (2Wiki) [17] subsets.
-Then, we establish the Closed-Book QA (CBQA) baseline for pre-trained Llama2-7B, which aims to identify the knowledge that has already been equipped from the newly injected knowledge.
-Consequently, we follow the traditional CPT+SFT paradigm to inject passage contents into model parameters, and supplement the experiment of our SCPT+SSFT technique for comparison.
-According to the results shown in Tab.
-1, the traditional continual pre-training strategy immediately improves the QA performance on six out of seven subsets (except the Qasper [10] subset), while the supervised fine-tuning alignment does not bring significant gains on the overall F1-Score measure.
-In contrast, our SCPT+SSFT approach successfully boosts the closed-book QA performance to 20.1% on average, even surpassing the open-book QA baseline of 18.7%.
-It indicates the vanilla SFT strategy can only regularize LLMs’ response styles, while our SSFT could teach LLMs to utilize their knowledge (injected in the SCPT stage) to answer corresponding questions, as exemplified in Fig.
+First, we use a CoT instruction of “Answer the question and explain why” to set a baseline for Open-Book QA (OBQA).
+As shown in Tab.
+1, it presents a moderate knowledge comprehending capability in SingleDoc QA tasks (e.g., over 40% recall on Qasper and MFQA), and relatively poor knowledge coordination in MultiDoc QA subsets (e.g., around 10% recall on Musiq and Duzh) due to the attention drift on longer inputs.
+Then, we try to inject passage content into LLMs with a conventional CPT+SFT paradigm to benchmark the Closed-Book QA (CBQA) baseline.
+Note that we use two types of question templates (refer to Sec.
+3.3) during training, and use the CoT version during testing to elicit models’ memorized knowledge in their responses.
+However, Tab.
+1 indicates such a knowledge injection approach is ineffective, as the knowledge recall is even lower than the OBQA baseline (i.e., 24.2% v.s.
+26.9%).
+On the contrary, our SCPT strategy has already achieved a considerable knowledge recall of 26.7%, which is higher than the CPT-SFT paradigm of 24.2% and approaching the OBQA baseline of 26.9%.
+It implies our structure-aware continually pre-trained model has successfully associated the relevant passages with their entire knowledge structure for the given question, which provides the knowledge path to seek targeted information to derive the answer.
+Furthermore, our SSFT technique explicitly teaches the model to recall the learned knowledge and answer the questions, which further improves the knowledge recall to 31.5% and even surpasses the Open-Book QA setting.
+The results indicate the vanilla SFT strategy can only regularize LLMs’ response styles, while our SSFT could teach LLMs to utilize their knowledge (injected in the SCPT stage) to answer corresponding questions, as exemplified in Fig.
 6.
-Table 1: F1 Score evaluation of Open-Book QA (OBQA) and Closed-Book QA (CBQA) tasks on the LongBench [4] dataset.
-The best results are marked in bold, and the secondary results are marked with underlines.
-The backbone model is Llama2-7B [40].
-Task Adaptation SingleDoc-QA MultiDoc-QA Average
-Qasper MFQA MFQAzh HpQA 2Wiki Musiq Duzh
-OBQA - 19.2 36.8 11.9 25.4 32.8 9.4 5.2 18.7
-CBQA - 17.3 15.9 1.4 5.7 6.7 3.3 1.8 7.4
-CPT 16.8 23.6 17.5 18.0 18.6 6.7 14.3 16.5
-CPT+SFT 16.8 23.1 13.2 21.3 19.1 10.4 13.4 16.8
-SCPT 15.2 21.5 15.2 14.9 19.8 5.6 14.1 15.2
-SCPT+SSFT 19.7 23.5 19.5 26.4 24.1 12.2 15.4 20.1
-Table 2: Recall evaluation of the Closed-Book QA task. The best results are marked in bold.
-Adaptation SingleDoc-QA MultiDoc-QA Average
-Qasper MFQA MFQAzh HpQA 2Wiki Musiq Duzh
-CPT 13.4 33.0 16.6 28.1 27.7 15.9 13.1 21.1
-CPT+SFT 20.7 35.3 20.6 29.9 32.1 18.9 12.0 24.2
-SCPT 18.8 42.5 17.7 35.7 36.4 20.5 15.3 26.7
-SCPT+SSFT 30.5 44.6 24.3 40.8 42.0 21.8 16.8 31.5
-To further support this claim, we compute Recall between model outputs and ground-truth answers in Tab.
-2 as a proxy measure for memorizing and manipulating the injected knowledge [59].
-Specifically, we use a CoT instruction of Answer the question and explain why.
-to elicit models’ knowledge in their responses.
-As shown in Tab.
-2, our SCPT strategy has already achieved higher knowledge recall than the CPT-SFT paradigm (26.7% v.s.
-24.2%).
-It implies our structure-aware continually pre-trained model has successfully associated the relevant passages with their entire knowledge structure for the given question, which provides the knowledge path to seek targeted information to derive the answer.
-As displayed in Fig.
-6, our SSFT technique explicitly teaches the model to recall the learned knowledge and answer the questions, which further improves the knowledge recall in Tab.
-2.
+Particularly, our method receives significant enhancements on MultiDoc QA subsets, which implies even though the total text content may exceed the models’ attention window and influence the OpenBook-QA, we can still chunk the content into several pieces and successfully inject the knowledge points into LLMs meanwhile preserving the whole knowledge structure.
 Figure 1: MindMap Recall
@@
 0.61 0.87
-In addition, we also statistically quantify the memorization of injected knowledge structures by comparing the mindmap in models’ responses with the ground-truth answers via lexical ROUGE-L [27] and semantic BERTScore [54] measures.
-The results shown in Fig.
-7 indicate a relatively good memorization of the injected knowledge mindmap, showing the efficacy of our SCPT strategy.
+In addition, we also use lexical ROUGE-L (Lin, 2004) and semantic BERTScore (Zhang et al., 2020) to quantify the memorization of injected knowledge structures, by comparing the mindmap in models’ responses (as Fig.
+6 displays) with ground-truth answers.
+The results in Fig.
+1 indicate a relatively good memorization of the injected knowledge mindmap, emphasizing the efficacy of our SCPT strategy.
 4.2 In-depth Evaluation for Multi-choice QA Application
 Datasets and Tasks.
-MMedBench [35] is a multilingual medical multi-choice QA benchmark, with 45,048 QA pairs for adapting LLMs to medical experts and 8,518 for testing.
-We collect 76M textbook corpora from MedTextBooks [21] and MMedC [35] for medical knowledge injection and use the training/test split from MMedBench for SFT/evaluation.
+MMedBench (Qiu et al., 2024) is a multilingual medical multi-choice QA benchmark, with 45,048 QA pairs for adapting LLMs to medical experts and 8,518 for testing.
+We collect 76M textbook corpora from MedTextBooks (Jin et al., 2020) and MMedC (Qiu et al., 2024) for medical knowledge injection and use the training/test split from MMedBench for SFT/evaluation.
+We also curate another two sizes of training sets (with 30M and 132M tokens) to validate our method’s scalability.
 Detailed setup is in Sec.
 A.2.
-Evaluation Metrics. For the multi-choice QA task in MMedBench [35], we follow the default setting to calculate the accuracy on six language subsets, as well as the averaged scores.
+Evaluation Metrics.
+For multi-choice QA in MMedBench (Qiu et al., 2024), we follow the default setting to calculate the accuracy on six language subsets, as well as the averaged scores.
+Metrics are computed by lexical exact-matching on models’ responses, rather than maximum token probabilities.
 Investigated Models.
-For MMedBench, we extend the investigated LLMs across model scales and architectures including Llama2-7B, Llama2-13B [40], and InternLM2-7B [57].
-We also compare our knowledge-injected models with other open- or closed-source LLMs, such as the domain-specified MedAlpaca [15], ChatDoctor [53], PMC-LLaMA [47], and MMedLM [35] models.
+For MMedBench, we extend the investigated LLMs across model scales and architectures including Llama2-7B/13B (Touvron et al., 2023b), InternLM2-7B (Zheng et al., 2024), and the recent Llama3-8B (Dubey et al., 2024).
+We also compare our knowledge-injected models with other popular domain-specified LLMs, such as MedAlpaca (Han et al., 2023), ChatDoctor (Yunxiang et al., 2023), PMC-LLaMA (Wu et al., 2024), and MMedLM (Qiu et al., 2024) models.
 Implementation Details.
@@
 In the SFT phase, the learning rate is set as 1e-6 to avoid overfitting on such an amount of SFT samples, as suggested by Qiu et al.
-[35].
+(2024).
 Main Results.
 In Tab.
-3, we present the overall performance across a series of LLMs on the testing split of MMedBench.
-The results demonstrate the promising enhancement achieved by our StructTuning technique, which translates into consistently significant improvements on Llama2-7B (+8.78%), Llama2-13B (+6.17%), and InternLM2-7B (+4.46%).
-It shows the generalizability and scalability of our StructTuning strategy across model architectures and sizes.
-Notably, our method, comprising SCPT on 76M curated textbook corpora and SSFT on 78K QA samples, achieves 50% performance (4.46% v.s.
-8.76%) against the state-of-the-art MMedLM2-7B model, which is trained on a huge MMedC [35] corpora of 25.5B tokens.
-Our approach shows much more efficiency in transforming pre-trained LLMs into domain experts by structured-aware knowledge injection, which largely reduces the cost for LLMs’ domain adaptation to 0.3%.
-However, our method is currently unable to achieve 100% of knowledge injection efficacy against MMedLM2, which may comprise two major reasons: (1) some of the knowledge in the training corpus has already been learned by the foundation InternLM2 model, and (2) the knowledge in our curated 76M training data cannot cover all the testing samples.
-We leave the promotion of achieving 100% improvement in our future work.
-Table 3: Multiple-choice accuracy evaluation on MMedBench [35].
+2, we present the overall performance across a series of LLMs on the testing split of MMedBench.
+The results demonstrate the promising enhancement achieved by our StructTuning technique, which translates into consistent improvements on the advanced InternLM2-7B (Zheng et al., 2024) and Llama3-8B (Dubey et al., 2024) models, and largely outperforms the previous domain-specific LLMs like PMC-LLaMA (Wu et al., 2024) and MedAlpaca (Han et al., 2023).
+Notably, our structure-aware knowledge injection approach, using merely 76M tokens curated from medical textbooks, achieves over 50% performance (4.46% v.s.
+8.71%, 2.57% v.s.
+4.96%) against the state-of-the-art MMedLM (Qiu et al., 2024) method, which is trained on a huge MMedC (Qiu et al., 2024) corpora of 25.5B tokens.
+Our approach shows much more efficiency in transforming pre-trained LLMs into domain experts by structured-aware knowledge injection, e.g., we only use 0.3% corpus to achieve comparable or even slightly better performance on English and Spanish subsets on top of Llama3-8B.
+However, our method is currently unable to achieve 100% of knowledge injection efficacy against MMedLM, which may comprise two major reasons: (1) some knowledge in the training corpus has already been learned by the foundation model (e.g., a slight decrease in the Chinese subset), and (2) knowledge in the curated 76M training data cannot cover all testing scenarios (e.g., a considerable gap on the Japanese subset due to the absence of Japanese corpus, see Tab.
+A2).
+Therefore, we make a step to evaluate our approach’s scalability on various corpora sizes in Fig.
+8.
+Table 2: Multiple-choice accuracy evaluation on MMedBench (Qiu et al., 2024).
 We report each model’s accuracy across six languages separately, with “Average” denoting the mean score over six languages.
-We also list out the data quantity for CPT and SFT stages during medical knowledge injection.
-Model CPT SFT English Chinese Japanese French Russian Spanish Average
-InternLM-7B - 45K 44.07 64.62 37.19 24.92 58.20 44.97 45.67 +0.00
-MMedLM-7B 25B 45K 49.88 70.49 46.23 36.66 72.27 54.52 55.01 +9.34
-InternLM2-7B - 45K 57.27 77.55 47.74 41.00 68.36 59.59 58.59 +0.00
-MMedLM2-7B 25B 45K 61.74 80.01 61.81 52.09 80.47 67.65 67.30 +8.71
-Llama2-7B - 45K 43.36 50.29 25.13 20.90 66.80 47.10 42.26 +0.00
-+Ours 76M 78K 49.41 65.15 36.68 35.21 69.14 50.62 51.04 +8.78
-Llama2-13B - 45K 51.37 57.97 32.66 25.08 69.92 52.99 48.33 +0.00
-+Ours 76M 78K 53.02 68.30 37.78 41.71 70.70 55.51 54.50 +6.17
-InternLM2-7B - 45K 57.27 77.55 47.74 41.00 68.36 59.59 58.59 +0.00
-+Ours 76M 78K 60.80 79.19 50.75 45.34 75.39 66.85 63.05 +4.46
+We also compare the data requirement (“#Token”) for medical knowledge injection.
+Model English Chinese Japanese French Russian Spanish Average #Token
+ChatDoctor 43.52 43.26 25.63 18.81 62.50 43.44 39.53 -
+PMC-LLaMA 47.53 42.44 24.12 20.74 62.11 43.29 40.04 -
+MedAlpaca 46.74 44.80 29.64 21.06 59.38 45.00 41.11 -
+Llama2-7B 43.36 50.29 25.13 20.90 66.80 47.10 42.26 -
+InternLM-7B 44.07 64.62 37.19 24.92 58.20 44.97 45.67 -
+InternLM2-7B 57.27 77.55 47.74 41.00 68.36 59.59 58.59 +0.00 -
++MMed 61.74 80.01 61.81 52.09 80.47 67.65 67.30 +8.71 25.5B
++Ours 60.80 79.19 50.75 45.34 75.39 66.85 63.05 +4.46 76M
+Llama3-8B 63.86 78.23 48.24 50.80 71.48 64.15 62.79 +0.00 -
++MMed 66.06 79.25 61.81 55.63 75.39 68.38 67.75 +4.96 25.5B
++Ours 66.77 77.44 53.27 51.61 74.61 68.49 65.36 +2.57 76M
+Figure 8: Comparison of knowledge injection approaches.
+Approach’s Scalability.
+In Tab.
+2, we use roughly 0.3% of the 25.5B tokens in MMedC Qiu et al.
+(2024) to evaluate knowledge injection methods.
+Here, another two training corpora sizes are curated: 30M and 132M, which take around 0.1% and 0.5% of 25.5B tokens respectively.
+We compare the vanilla CPT-SFT paradigm and our SCPT-SSFT strategy across those data settings The backbone LLM is InternLM2-7B (Zheng et al., 2024), and the training settings follow the main experiments.
+According to Fig.
+8, our method consistently surpasses the vanilla paradigm by a large margin, emphasizing the efficacy and efficiency of domain knowledge injection.
+In particular, we can fit two performance-ratio scaling curves from the data points in Fig.
+8 as:
+$p_{v}\approx-0.04(\log r)^{2}+13.3\log r+100.0;\quad p_{s}\approx-1.11(\log r)^{2}+7.63\log r+133.0$ (3)
+where $p_{v}$ and $p_{s}$ denote the relative performance enhancement (%) for vanilla and structure-aware knowledge injection, and $r$ refers to the corpus ratio.
+The scaling law indicates we can achieve comparable performance (100%) with only 5% of the total training corpus, significantly reducing the costs for LLMs’ domain adaptation.
+On the other hand, it also indicates our method may lead to 133% enhancement with a further 100% comprehensive data utilization.
+However, we have not empirically verified those predictions due to time and resource constraints, which we view as a temporal limitation of our current work and leave it to future investigations.
+Approach’s Generalization.
+In addition to the InternLM2 and Llama3 models in previous experiments, we also investigate the generalization ability of our SCPT+SSFT paradigm on Llama2 (Touvron et al., 2023b) model series.
+As shown in Tab.
+3, our method leads to consistently significant improvements on 7B (+8.78%) and 13B (+6.17%) backbone models.
+The results further demonstrate the generalizability and scalability of our StructTuning strategy across model architectures and sizes.
+Table 3: Structure-aware knowledge injection to Llama2 (Touvron et al., 2023b) model series.
+Model English Chinese Japanese French Russian Spanish Average
+Llama2-7B 43.36 50.29 25.13 20.90 66.80 47.10 42.26
++Ours 49.41 65.15 36.68 35.21 69.14 50.62 51.04
+Llama2-13B 51.37 57.97 32.66 25.08 69.92 52.99 48.33
++Ours 53.02 68.30 37.78 41.71 70.70 55.51 54.50
 Ablation Studies.
-We conduct a comprehensive ablation study with the English split of the MMedBench dataset to evaluate our efficacy.
-Specifically, we take Llama2-7B as the backbone model, select the English textbooks [21] (with 26M tokens) for vanilla and our structure-aware continual pre-training, and use different QA samples for supervised fine-tuning.
-In particular, in Tab.
-4, SFT refers to vanilla SFT with 10K English QA samples provided in MMedBench’s [35] training split, SSFT indicates structure-aware SFT on 10K training data, where the questions are the same as SFT while the answers are enhanced with knowledge structure explanation by Llama3-70B, as described in Sec.
+To further demonstrate the efficacy of each component of our StructTuning method, we conduct a comprehensive ablation study with the English split of the MMedBench dataset.
+Specifically, we take Llama2-7B as the backbone model, select the English textbooks (Jin et al., 2020) (with 26M tokens) for vanilla and our structure-aware continual pre-training, and use different QA samples for supervised fine-tuning.
+In Tab.
+4, “SFT” refers to vanilla SFT with 10K English QA samples provided in MMedBench’s (Qiu et al., 2024) training split, “SSFT” indicates structure-aware SFT on 10K training data, where the questions are the same as “SFT” while the answers are enhanced with knowledge structure explanation by Llama3-70B, as described in Sec.
 3.3.
-SSFT* includes another 8K structure-aware QA syntheses by Llama3-70B, consisting of 18K entries for training.
+“SSFT*” includes another 8K structure-aware QA syntheses by Llama3-70B, consisting of 18K entries for training.
 The training hyper-parameters follow the main experiment.
-Results in Tab.
-4 imply the vanilla CPT+SFT injection paradigm does bring considerable improvement of 1.73% in the English split, while our SCPT technique with vanilla SFT strategy presents a slightly higher accuracy of 46.50%.
-However, when combining SSFT with SCPT, our method immediately leads to a significant boost in the English split (49.96% v.s.
-44.54%), demonstrating the efficacy and necessity of eliciting the learned structured knowledge to solve practical problems.
-Moreover, the supplementation of 8K extra QA syntheses surprisingly enhances the model performance on the other five subsets, resulting in a valuable gain on the average accuracy of 38.27%.
-This result demonstrates the knowledge transferability across different languages [24, 34].
-After training with our SSFT strategy, LLMs can actively utilize the knowledge injected in one language to solve the problem in another language, further evidencing our superiority against the traditional SFT technique.
-In addition, we observe that the commonly used RAG [25] strategy does not bring significant advantages to the MMedBench evaluation.
-The main reason lies in the gap between the pre-training corpus (comprising official knowledge statements from textbooks) and evaluated QA samples (originating from practical diagnosis records).
-To teach LLMs to apply relevant medical knowledge for reliable diagnosis answers, knowledge injection by CPT and SFT shows more advantages in this situation.
-Table 4: Ablation studies with Llama2-7B on the English subset of MMedBench. The best results are marked in bold, and the secondary results are marked with underlines.
+Table 4: Ablation studies with Llama2-7B on the English subset of MMedBench.
 Adaptation English Chinese Japanese French Russian Spanish Average
@@
 RAG 38.12 29.22 22.61 23.34 53.91 36.47 33.95
+In Tab.
+4, vanilla CPT+SFT injection paradigm does bring an improvement of 1.73% in the English split, while our SCPT technique with vanilla SFT strategy presents a slightly higher accuracy of 46.50%.
+However, when combining SSFT with SCPT, our method immediately leads to a significant boost in the English split (49.96% v.s.
+44.54%), demonstrating the efficacy and necessity of eliciting the learned structured knowledge to solve practical problems.
+Moreover, the supplementation of 8K extra QA syntheses (“SSFT*”) surprisingly enhances the model performance on the other five subsets, resulting in a valuable gain on the average accuracy to 38.27%.
+This result demonstrates the knowledge transferability across different languages (Lai et al., 2023; Qin et al., 2024).
+After training with SSFT, LLMs can actively utilize the knowledge injected in one language to solve the problem in another language, evidencing our superiority against the traditional SFT technique.
+In Sec.
+B.4, we follow Liu et al.
+(2024b) to randomly generate another 8K SFT pairs for further comparison, where our structure-aware SFT syntheses are verified to better enhance LLMs’ knowledge application.
+In addition, we observe that the commonly used RAG (Lewis et al., 2020) strategy does not bring significant advantages to the MMedBench evaluation.
+The main reason lies in the gap between the pre-training corpus (comprising official knowledge statements from textbooks) and evaluated QA samples (originating from practical diagnosis records).
+Knowledge injection by (S)CPT and (S)SFT shows more advantages in this situation.
 5 Conclusion
-This work pioneers in incorporating structure-aware methodologies to enhance domain knowledge injection into large language models (LLMs).
-Through a novel two-stage training strategy, combining the SCPT and SSFT techniques, we have set a new precedent for efficiently adapting LLMs to specialized domains.
-The promising results and the scalable data construction framework we developed underscore the viability and potential of our method.
-The limitations and broader impacts are discussed in Appendix C.
+This work pioneers in incorporating structure-aware methodologies to enhance domain knowledge injection into large language models.
+Through a novel SCPT-SSFT paradigm, we have set a new precedent for adapting LLMs to specialized domains, and the promising and scalable results underscore the viability and potential of our method.
 We hope to inspire further research in efficient and effective domain adaptation, moving a step closer to models that can truly emulate human intelligence.
+Limitation.
+A noteworthy limitation of our work is the under-explored full potential of our StructTuning strategy for knowledge injection.
+Although we have estimated a scaling law in Eq.
+3, the data requirement to achieve 100% of the desired effectiveness remains unverified.
+Besides, our method introduces negligible data preprocessing costs to extract the domain knowledge structure.
+However, Sec.
+B.2 indicates our method can still reduce the overall adaptation cost to 10% since only a small part of training corpus is required.
+We will delve into the investigations in our future work.
 Appendix A Implementation Details
 A.1 Detailed Setup on LongBench
-To focus on the investigation of knowledge injection, we choose 7 subsets from LongBench [4] across single- and multi-document QA tasks in English and Chinese, and the remaining synthetic or code-orientated tasks are eliminated:
+To focus on the investigation of knowledge injection, we choose 7 subsets from LongBench (Bai et al., 2023) across single- and multi-document QA tasks in English and Chinese, and the remaining synthetic or code-orientated tasks are eliminated:
   • Single-Doc QA.
-For single-document QA, we take three subsets from LongBench: (1) Qasper [10], featured by question-answering over NLP technical papers and annotated by NLP practitioners; (2) MultiFieldQA [4], manually curated from multiple data sources and annotated by Ph.D.
+For single-document QA, we take three subsets from LongBench: (1) Qasper (Dasigi et al., 2021), featured by question-answering over NLP technical papers and annotated by NLP practitioners; (2) MultiFieldQA (Bai et al., 2023), manually curated from multiple data sources and annotated by Ph.D.
 students; and (3) MultiFieldQA-zh, the Chinese split also provided by Bai et al.
-[4], covering multiple Chinese scenarios.
+(2023), covering multiple Chinese scenarios.
 MultiFieldQA contains 150 Context-Question-Answer triplets to test, and the others adopted subsets include 200 pieces of test samples respectively.
@@
 Multi-document QA requires LLMs to extract and combine information from multiple documents to derive the answer, which is generally more challenging than single-doc QA.
-We take four multi-hop QA datasets: (1) HotpotQA [51], containing 2-hop questions written by native speakers given two related paragraphs; (2) 2WikiMultihopQA [17], involving up to 5-hop questions synthesized through manually designed templates on Wikipedia passages; (3) MuSiQue [41], carefully composed with up to 4-hop reasoning on an increased number of supporting and distracting context evidence; and (4) Dureader [16], developed based on Baidu Search and Baidu Zhidao and filtered by Bai et al.
-[4] to reduce the data noise.
+We take four multi-hop QA datasets: (1) HotpotQA (Yang et al., 2018), containing 2-hop questions written by native speakers given two related paragraphs; (2) 2WikiMultihopQA (Ho et al., 2020), involving up to 5-hop questions synthesized through manually designed templates on Wikipedia passages; (3) MuSiQue (Trivedi et al., 2022), carefully composed with up to 4-hop reasoning on an increased number of supporting and distracting context evidence; and (4) Dureader (He et al., 2017), developed based on Baidu Search and Baidu Zhidao and filtered by Bai et al.
+(2023) to reduce the data noise.
 Each subset has 200 test samples.
+For data entries in Single-Doc QA, we extract the knowledge structure for each single passage; in Multi-Doc QA, we identify the knowledge structure across multiple passages for each test sample.
+There are ultimate 1350 question-answer-passage(s)-(knowledge)structure quadruples to evaluate knowledge injection approaches on LongBench.
 A.2 Detailed Setup on MMedBench
 Data for Evaluation.
-The Multilingual Medical Benchmark (MMedBench) [35] represents a comprehensive and diverse multilingual medical Question and Answering (QA) benchmark designed to evaluate models’ capabilities of understanding and processing medical content.
+The Multilingual Medical Benchmark (MMedBench) (Qiu et al., 2024) represents a comprehensive and diverse multilingual medical Question and Answering (QA) benchmark designed to evaluate models’ capabilities of understanding and processing medical content.
 MMedBench’s robust dataset extends across 6 languages (i.e., English, Chinese, Japanese, French, Russian, and Spanish) and 21 medical fields, which include, but are not limited to, Internal Medicine, Biochemistry, Pharmacology, Psychiatry, and many others.
@@
 Data for Continual Pre-Training.
-To investigate high-quality domain knowledge injection for LLMs, we collect 18 English textbooks and 33 Chinese textbooks from the National Medical Board Examination in the USA and Mainland China, respectively [21].
+To investigate high-quality domain knowledge injection for LLMs, we collect 18 English textbooks and 33 Chinese textbooks from the National Medical Board Examination in the USA and Mainland China, respectively (Jin et al., 2020).
 All collected textbooks are originally in PDF format and Jin et al.
-[21] converted them into digital text via OCR and performed some clean-up pre-processing strategies to reduce the data noise.
-The English and Chinese textbooks count for around 26.1M and 21.5M tokens by Llama2 tokenizer [40].
-Then, we randomly sample an extra 28M textbook corpora from MMedC [35] for the other languages (except the unavailable Japanese textbooks), and the final statistics are displayed in Tab.
+(2020) converted them into digital text via OCR and performed some clean-up pre-processing strategies to reduce the data noise.
+The English and Chinese textbooks count for around 26.1M and 21.5M tokens by Llama2 tokenizer (Touvron et al., 2023b).
+Then, we randomly sample an extra 28M textbook corpora from MMedC (Qiu et al., 2024) for the other languages (except the unavailable Japanese textbooks), and the statistics are displayed in Tab.
 A2.
-Table A2: Sample statistics on training data.
+Table A2: Sample statistics on training data. $\dagger$” and $\ddagger$” refer to different sizes.
 Stage English Chinese Japanese French Russian Spanish Total
 CPT 26.1M 21.5M - 8.1M 10.3M 10.1M 76.1M
+[…]
+[…]
 SFT 18.8K 39.1K 1.6K 5.3K 5.9K 7.5K 78.2K
+[…]
 Data for Supervised Fine-Tuning.
 As introduced in Sec.
-3.3, we prompt Llama3-70B [2] to build the structure-aware answer explanations on top of the raw SFT samples in MMedBench’s training split, and generate extra QA pairs by traversing the extracted knowledge structure from textbooks.
+3.3, we prompt Llama3-70B (Dubey et al., 2024) to build the structure-aware answer explanations on top of the raw SFT samples in MMedBench’s training split, and generate extra QA pairs by traversing the extracted knowledge structure from textbooks.
 The final quantity statistics are presented in Tab.
 A2.
-Appendix B Resource Requirement
+Knowledge Structures.
+We extract the domain knowledge structure for each textbook, where Fig.
+3 presents an example, and combine the medical knowledge for six languages in MMedBench.
+As the (S)CPT corpus for Japanese is collected from Wikipedia rather than textbooks, we derive a single knowledge structure for Japanese medicine.
+Figure A1: Knowledge structure extraction from (a) sequential chunks (e.g., from textbooks) by our specialized 7B model and (b) separated trunks (e.g., from websites) by clustering-based methods (Sarthi et al., 2024).
+Here the terms of “Section”, “Chapter”, and “book” are just examples to help illustrate the knowledge structure.
+A.3 Resource Requirement
 We use 8 NVIDIA A100-80G GPUs to train all the models, and leverage 1-2 NVIDIA A100-80G GPUs for inference.
-Appendix C Limitations and Discussions
-Limitations.
-A noteworthy limitation of our work, despite its significant achievements, is centered around the exploration of the full potential of our StructTuning strategy for knowledge injection.
-While our method has successfully demonstrated a remarkable 50% improvement in domain-specific knowledge application and retention using merely 0.3% of the training corpus, the extent of data scalability needed to achieve 100% of the desired effectiveness remains unexplored.
-This limitation primarily arises from constraints in time and resources available during our research phase.
-We will delve into the investigations in our future work.
-Broader Impacts. We discuss the positive and negative societal impacts as follows:
-  • Positive Societal Impacts.
-Our StructTuning strategy could democratize access to advanced AI, enabling specialized AI assistants across various sectors such as healthcare and education.
-This method’s efficiency could improve diagnostic tools, personalized learning experiences, and enhanced decision-making processes.
-Furthermore, by mirroring the human learning process, our approach facilitates the development of AI systems that are easier for people to trust and interact with.
-  • Negative Societal Impacts.
-The enhanced capabilities of LLMs might be exploited for generating convincing yet fraudulent content, raising concerns over misinformation and intellectual deceit.
-Additionally, if the structured knowledge base contains biases, it could inadvertently reinforce societal stereotypes, necessitating rigorous curation to ensure fairness and accuracy.
-To counteract potential misuse, controlled model releases and continuous monitoring for abuse are recommended.
-Ensuring the diversity and contemporaneity of training data can mitigate bias.
-Collaborative efforts with domain experts, ethicists, and policymakers are vital in developing guidelines for the technology’s ethical use.
-Appendix D Prompt Template for Structurization
+Appendix B Additional Experiments
+B.1 Ablation on Knowledge Structure Extraction
+Extracting domain knowledge structure is a prerequisite for subsequent knowledge injection (including both SCPT and SSFT) for language models.
+In Sec.
+3.1, we propose a bottom-up strategy to re-chunk the texts from domain textbooks, summarize a title for each chunk, and send the title list to a specialized 7B model to derive the knowledge structure.
+The prompt template is displayed in Appendix C.
+In fact, we argue that due to the language diversity, a perfectly recovered table of contents of textbooks is unnecessary for domain knowledge injection.
+A reasonable knowledge structure is sufficient enough.
+In Tab.
+A3, we individually adopt few-shot GPT-3.5-Turbo (Brown et al., 2020) and LLaMA3-70B (Dubey et al., 2024) models to extract medical knowledge structure from 18 English textbooks (Jin et al., 2020) (with 26.1M tokens) for subsequent knowledge injection (the backbone LLM is LLaMA2-7B (Touvron et al., 2023b)).
+Although they present 3.6%-3.7% enhancement on MMedBench (Qiu et al., 2024)’s English test set (denoted as “improvement”), leveraging GPT-3.5-Turbo and LLaMA3-70B is either expensive or time-consuming.
+GPT-3.5-Turbo costs around 15 dollars to process 26M tokens, while LLaMA3-70B takes around 1.5 hours on 2 A100-80G GPUs, of which both limit the scalability of data pre-processing for structure-aware knowledge injection.
+Table A3: Comparison of models to extract knowledge structures on 26M English corpus.
+Model Improvement Time Consumption Extra Cost
+GPT-3.5-Turbo +3.58 - 15\$
+LLaMA3-70B +3.72 1.5h -
+Ours-7B +3.69 0.2h -
+Inspired by Liu et al.
+(2024a), we distilled the knowledge structure extraction capability from giant LLMs to a LLaMA2-7B model via supervised fine-tuning.
+In particular, we instruct LLaMA3-70B to generate 22K training examples (pairs of raw knowledge points and extracted knowledge structures) from Wikipedia, and train a LLaMA2-7B model at a batch size of 128 and a learning rate of 2e-5 for 1 epoch.
+After utilizing the specialized 7B model to identify the knowledge structure in medical textbooks, as shown in Tab.
+A3, the results translate to comparable performance on structure-aware knowledge injection.
+Meanwhile, the inference cost significantly decreases to 0.3 hours, which is more scalable to handle a larger amount of domain corpus.
+B.2 Comparison on Training Costs
+In Tab.
+A4, we quantify the total training cost on 8 A100-80G GPUs.
+According to Qiu et al.
+(2024), the conventional CPT+SFT paradigm on 25.5B medicine corpus takes more than 30 days to derive the SOTA MMedLM model.
+In our SCPT+SSFT framework, although the pre-processing (i.e., knowledge structure extraction) introduces an extra 0.6 hours to process 0.3% data (around 76M tokens), the total training process only costs 4.5 hours.
+As suggested in Fig.
+8, when 5% training data is leveraged for knowledge injection to achieve 100% improvement, the overall cost is limited to 3 days, much less than the CPT+SFT approach with more than a month.
+Those analyses further demonstrate the efficacy and efficiency of our structure-aware knowledge injection framework.
+Table A4: Comparison of training costs for knowledge injection.
+Paradigm Corpus Improvement Pre-process Total
+CPT+SFT 100% 100% - $>$30d
+SCPT+SSFT 0.3% 50% 0.6h 4.5h
+SCPT+SSFT 5% 100% 9.7h 3d
+B.3 F1-Score Evaluation on LongBench
+In Sec.
+4.1, we mainly follow Zhu & Li (2023b) to investigate the memorization and understanding of injected knowledge by calculating the knowledge recall in models’ responses.
+Here we report the F1-score measure over the Open-Book QA (OBQA) and Closed-Book QA (CBQA) settings for a thorough comparison.
+Note that here we use the vanilla question prompt to obtain concise answers, instead of the CoT prompt used in Sec.
+4.1 to elicit models’ memorized knowledge.
+The evaluated models are the same as Sec.
+4.1.
+In Tab.
+A5, we report the Open-Book QA (OBQA) baseline for Llama2-7B with passages as inputs, which shows the best performance on MultiFieldQA (MFQA) (Bai et al., 2023) and 2WikiMultihopQA (2Wiki) (Ho et al., 2020) subsets.
+Then, we establish the Closed-Book QA (CBQA) baseline by traditional CPT+SFT to inject passage contents into model parameters, and supplement the experiment of our SCPT+SSFT technique for comparison.
+According to the results shown in Tab.
+A5, CPT+SFT slightly improves the QA performance on several subsets (such as MultiFieldQA-zh (MFQAzh) (Bai et al., 2023)), while the overall F1-Score measure is still inferior to the OBQA performance.
+In contrast, our SCPT+SSFT approach successfully boosts the closed-book QA performance to 20.1% on average, even surpassing the open-book QA baseline of 18.7%.
+The results are consistent with Sec. 4.1, which jointly demonstrate the effectiveness of structure-aware knowledge injection for large language models.
+Table A5: F1 Score evaluation of Open-Book QA (OBQA) and Closed-Book QA (CBQA) tasks on the LongBench (Bai et al., 2023) dataset.
+The best results are marked in bold, and the secondary results are marked with underlines.
+The backbone model is Llama2-7B (Touvron et al., 2023b).
+Task Adaptation SingleDoc-QA MultiDoc-QA Average
+Qasper MFQA MFQAzh HpQA 2Wiki Musiq Duzh
+OBQA - 19.2 36.8 11.9 25.4 32.8 9.4 5.2 18.7
+CBQA CPT+SFT 16.8 23.1 13.2 21.3 19.1 10.4 13.4 16.8
+SCPT+SFT 15.2 21.5 15.2 14.9 19.8 5.6 14.1 15.2
+SCPT+SSFT 19.7 23.5 19.5 26.4 24.1 12.2 15.4 20.1
+B.4 Ablation on SFT Data Synthesis
+In Sec.
+4.2, we compared our structure-aware knowledge injection with conventional CPT+SFT paradigm on MMedBench.
+On its English subset, we ablated the training components of our method, and found that the newly synthesized 8K SSFT data (by traversing the extracted knowledge structure) can inspire LLMs’ cross-language capability to apply the learned structured knowledge to solve practical diagnosis problems.
+Here, we follow Liu et al.
+(2024b) to randomly generate another 8K QA pairs for SFT alignment for further comparison, denoted as “SFT*”.
+We randomly sample medical texts and instruct Llama3-70B (Dubey et al., 2024) for data synthesis, without the knowledge structure provided.
+Tab.
+A6 indicates that “SFT*” brings slight enhancement to the English test subset, but the average accuracy drops to 34.51% instead.
+The results further demonstrate our method’s efficacy in the application of the injected, structured domain knowledge.
+Table A6: Comparison of SFT data synthesis strategies on MMedBench. The backbone LLM is the same Llama2-7B model after SCPT on English textbooks.
+SFT synthesis English Chinese Japanese French Russian Spanish Average
+- 46.50 32.14 20.10 18.17 53.91 39.97 35.13
+SFT* 47.13 32.49 16.58 16.72 51.95 42.16 34.51
+SSFT* 49.10 33.92 18.33 27.14 57.42 43.73 38.27
+Appendix C Prompt Template for Knowledge Structure Extraction
 Fig.
-A1 displays the few-shot prompt template to query giant commercial LLMs to execute structurization.
-We first introduce the instruction for structurization (including the task definition and output formats), then provide two representative examples to further illustrate the process.
-The first example in Fig.
-A2 tells the model to focus on the existing numerical or enumeration indicators to assist in constructing the aspect-description structure.
-The second example in Fig.
-A3 teaches the model to automatically summarize the aspects and attach their corresponding descriptions from the raw text sequences.
-Finally, in Fig.
-A1 we ask commercial LLMs to structurize the $input\_statement$ from user input as expected.
-Figure A1: Prompt template for few-shot structurization.
-Figure A2: The first example for few-shot structurization.
-Figure A3: The second example for few-shot structurization.
+A2 displays the prompt template to query our specialized 7B model to extract knowledge structure on given knowledge points, which introduces the task definition, detailed instruction, and output formats to illustrate the process.
+Figure A2: Prompt template for knowledge structure identification.

### v3

@@
 This paper introduces a pioneering methodology, termed StructTuning, to efficiently transform foundation Large Language Models (LLMs) into domain specialists.
-It significantly reduces the training corpus requirement to a mere 0.3%, while achieving an impressive 50% of traditional knowledge injection performance.
-Our method is inspired by the educational processes of human students, particularly how structured domain knowledge from textbooks is assimilated and subsequently applied to tackle real-world challenges through specific exercises.
-Based on this, we propose a novel two-stage strategy for knowledge injection and alignment: Structure-aware Continual Pre-Training (SCPT) and Structure-aware Supervised Fine-Tuning (SSFT).
+It significantly reduces the training corpus needs to a mere 5% while achieving an impressive 100% of traditional knowledge injection performance.
+Motivated by structured human education, we propose a novel two-stage strategy for knowledge injection and alignment: Structure-aware Continual Pre-Training (SCPT) and Structure-aware Supervised Fine-Tuning (SSFT).
 In the SCPT phase, we automatically extract the domain knowledge taxonomy and reorganize the training corpora, enabling LLMs to effectively link textual segments to targeted knowledge points within the taxonomy.
 In the SSFT phase, we explicitly prompt models to elucidate the underlying knowledge structure in their outputs, leveraging the structured domain insight to address practical problems.
-Our ultimate method has undergone extensive evaluations across model architectures and scales, using closed-book question-answering tasks on LongBench and MMedBench datasets.
-Remarkably, our method demonstrates the potential of comparable improvement against the state-of-the-art MMedLM2 on MMedBench, while significantly reducing the training costs to 5%.
-This breakthrough paves the way for scaling up our StructTuning for stronger domain-specific LLMs with comprehensive data utilization.
-Code is available at https://github.com/alibaba/struxgpt.
+Our ultimate method was extensively evaluated across model architectures and scales on LongBench and MMedBench datasets, demonstrating superior performance against other knowledge injection methods.
+We also explored our method’s scalability across different training corpus sizes, laying the foundation to enhance domain-specific LLMs with better data utilization.
+Code is available at: https://github.com/alibaba/struxgpt.
+Structure-aware Domain Knowledge Injection for Large Language Models
+Kai Liu1,2111Equal contribution.
+Work done during Kai Liu’s research internship at Alibaba Cloud.
+Email: kail@zju.edu.cn.,  Ze Chen2111Equal contribution.
+Work done during Kai Liu’s research internship at Alibaba Cloud.
+Email: kail@zju.edu.cn.,  Zhihang Fu2222Corresponding authors.
+Email: zhihang.fzh@alibaba-inc.com, 0006236@zju.edu.cn.,  Wei Zhang1,  Rongxin Jiang1, Fan Zhou1222Corresponding authors.
+Email: zhihang.fzh@alibaba-inc.com, 0006236@zju.edu.cn.,   Yaowu Chen1,  Yue Wu2,  Jieping Ye2 1Zhejiang University,   2Alibaba Cloud
 1 Introduction
+Figure 1: Discrepancy between human education and vanilla LLM adaptation.
+Human students learn structured knowledge through textbooks section by section, with particular exercises on related knowledge points.
+Traditional LLM adaptation continually pre-trains on data chunks from randomly concatenated text segments, with aimless supervised fine-tuning for conversation alignment.
+The inherent knowledge structure is ignored.
 Large language models (LLMs) have recently seen extensive deployment across various applications (Vaswani et al., 2017; Achiam et al., 2023; Jiang et al., 2023; Bi et al., 2024).
-When adapting foundational models (e.g., Llama series (Touvron et al., 2023a; b; Dubey et al., 2024)) for specialized AI assistants in distinct domains (Qiu et al., 2024; Guo et al., 2024), developers usually employ two techniques to enhance LLMs’ proficiency: retrieval-augmented generation (RAG) (Lewis et al., 2020) and domain knowledge injection (Gururangan et al., 2020).
+When adapting foundational models (e.g., Llama series (Touvron et al., 2023a, b; Dubey et al., 2024)) to specialized AI assistants in distinct domains, developers usually employ two techniques to enhance LLMs’ proficiency: retrieval-augmented generation (RAG) (Lewis et al., 2020) and domain knowledge injection (Gururangan et al., 2020).
 While RAG effectively utilizes an external knowledge base to augment information, the retrieval process’s inherent noise poses challenges to generating reliable responses, especially in scenarios requiring logical reasoning where there is a semantic gap between the user’s query and the knowledge base.(Zhang et al., 2023; Chen et al., 2023).
-Thus, another avenue of research focuses on injecting new knowledge into models via training techniques (Gu et al., 2021; Hu et al., 2021; Mecklenburg et al., 2024).
-Continual pre-training (Sun et al., 2020; Ibrahim et al., 2024) has been preferred for integrating new, domain-specific knowledge into existing LLMs (Cui et al., 2023; Wang et al., 2023b; Qiu et al., 2024).
-Nevertheless, it often entails resource-intensive auto-regressive training on billions of tokens from the internet to learn fragmented knowledge points, rather than absorbing structured knowledge from a few domain-specific textbooks (Jin et al., 2020).
-For example, MMedLM (Qiu et al., 2024) curates 25.5B tokens to derive a medical model, and DeepSeek-Coder (Guo et al., 2024) uses 2T tokens for coding adaptation.
-The common failure to learn effectively from limited textbook content has been attributed to insufficient data diversity (Zhu & Li, 2023a), which however violates the observation during the human education process in Fig.
+Thus, another avenue tries to inject new knowledge to LLMs via training techniques (Gu et al., 2021; Hu et al., 2021; Mecklenburg et al., 2024).
+Continual pre-training (Sun et al., 2020; Ibrahim et al., 2024) is widely used for injecting domain-specific knowledge (Cui et al., 2023; Wang et al., 2023b; Qiu et al., 2024).
+However, it often requires training on billions of internet tokens to capture fragmented knowledge, rather than leveraging a few structured textbooks (Jin et al., 2020).
+For instance, MMedLM (Qiu et al., 2024) uses 25.5B tokens for medical modeling, while DeepSeek-Coder (Guo et al., 2024) processes 2T tokens for coding adaptation.
+The limited ability to learn effectively from textbooks was attributed to insufficient data diversity (Zhu and Li, 2023a), which however violates the observation during the human education process in Fig.
 1: students gain knowledge by sequentially studying from textbooks, reviewing knowledge points and structures, and applying this knowledge through proper exercises.
-In this process, all the new data to learn are textbooks (structured content) and exercising examples (question-answering pairs), and students just adopt their world knowledge to memorize, understand, and apply the knowledge to become domain experts (Krathwohl, 2002; Yu et al., 2023).
-Figure 1: Discrepancy between human learning process and vanilla LLM adaptation paradigm.
-Human students learn structured knowledge through textbooks section by section, with particular exercises on related knowledge points.
-Traditional LLM adaptation involves continual pre-training on data chunks from randomly concatenated text segments, with aimless supervised fine-tuning for conversation alignment.
-The inherent property of structured knowledge is ignored.
-Inspired by this, we propose to inject the domain knowledge from textbooks into LLMs, as educating a human student, through a novel two-stage training strategy: Structure-aware Continual Pre-Training (SCPT) and Structure-aware Supervised Fine-Tuning (SSFT).
-In the SCPT stage, we argue that high-quality textbook data can adequately infuse the domain knowledge (Gunasekar et al., 2023), where the organization of training corpora is crucial.
-In conventional paradigms, as illustrated in Fig.
-1, text corpora are simply concatenated and divided into chunks of 2048 (Qiu et al., 2024) or 4096 (Guo et al., 2024), while the inherent structure of the texts (e.g., catalogs of textbooks) is disregarded.
-Instead, we propose an automatic approach to maintain each chunk’s knowledge structure.
-We view each chunk as a knowledge point, and employ advanced LLMs to efficiently extract domain knowledge taxonomy from the corpus, bypassing the need for manual annotation.
-Subsequently, LLMs are trained to predict the textual content (corresponding to the knowledge point) under the condition of the knowledge path within the domain structure, linking individual training chunks with the entire knowledge architecture.
-Finally, models are asked to memorize the knowledge structure to review the whole domain knowledge system.
+Here, all the new data to learn are textbooks (structured content) and exercising examples (question-answering pairs), and students just adopt their world knowledge to memorize, understand, and apply the knowledge to become domain experts (Krathwohl, 2002; Yu et al., 2023).
+As educating human students, we propose to inject structured domain knowledge into LLMs via two steps: Structure-aware Continual Pre-Training (SCPT) and Structure-aware Supervised Fine-Tuning (SSFT).
+In the SCPT stage, we argue that high-quality textbook data (as well as regular corpora from the Internet) can adequately infuse domain knowledge (Gunasekar et al., 2023), where the organization of training corpora is crucial.
+In conventional paradigms (Fig.
+1), text corpora are simply concatenated and divided into chunks of 2048 (Qiu et al., 2024) or 4096 (Guo et al., 2024) tokens, while the inherent semantic structure (e.g., catalogs of textbooks) is discarded.
+Instead, we view each chunk as a knowledge point and automatically extract domain knowledge taxonomy from the whole corpus.
+Subsequently, LLMs are trained to predict textual content (corresponding to a knowledge point) under the condition of the knowledge path within the domain structure, linking individual training chunks with the entire knowledge architecture.
+Finally, models have to memorize the entire structure to review the whole domain knowledge system.
 In the SSFT stage, the goal shifts from knowledge injection to enabling LLMs to recall and utilize their acquired knowledge to tackle real-world challenges.
 We explicitly elicit knowledge paths in LLMs’ responses, as a beacon for models to targeted information retrieval or logical reasoning for reliable responses.
-To this end, we derive a scalable strategy to generate question-answer pairs as practice exercises by powerful LLMs such as GPT4 (Achiam et al., 2023) or LLaMA3 (Dubey et al., 2024).
-In the scenarios with existing QA pairs like MMedBench (Qiu et al., 2024), we retrieve related knowledge structure and content, instructing LLaMA3 to provide explanations from questions to answers based on the knowledge paths.
-For datasets lacking specific QA samples like LongBench (Bai et al., 2023), we randomly select knowledge paths from the domain taxonomy and prompt LLaMA3 to craft question-answer-explanation triplets for training exercises.
-Our ultimate approach, termed StructTuning, outperforms conventional methods in domain knowledge injection by emulating human learning processes through SCPT and SSFT phases.
-We extensively evaluate StructTuning’s effectiveness across different model architectures and sizes.
-For domain-adapted language models, we first examine their capability to recall the injected knowledge through open-ended QA on the LongBench (Bai et al., 2023) dataset, then assess their application of this knowledge in addressing real-world issues via multiple-choice QA on MMedBench (Qiu et al., 2024).
-Both evaluations underscore the superiority of StructTuning.
-Remarkably, we achieve a 50% improvement in knowledge injection compared to the SOTA MMedLM2 in the medical domain, using merely 0.3% of the training data requirement.
-And StructTuning illustrates the potential of comparable performance with only 5% of training costs.
-These findings reveal the scalability of our method for enhancing domain-specific AI assistants with further comprehensive data utilization.
+To this end, we derive a scalable strategy to generate question-answer pairs as practice exercises by open-sourced LLMs or API, such as LLaMA3 (Dubey et al., 2024) and GPT4 (Achiam et al., 2023).
+In the scenarios with existing QA pairs like MMedBench (Qiu et al., 2024), we retrieve the related knowledge structure and content, instructing LLaMA3 to provide explanations from questions to answers based on the knowledge paths.
+For datasets lacking specific QA samples like LongBench (Bai et al., 2023b), we randomly select knowledge paths from the domain taxonomy and prompt LLaMA3 to craft question-answer-explanation triplets for training exercises.
+Our ultimate approach StructTuning has been extensively evaluated across different model architectures and sizes.
+In particular, we first examine their capability to recall the knowledge injected through open-ended QA on the LongBench (Bai et al., 2023b) dataset, then assess the application of injected knowledge to address real-world issues through multiple-choice QA on MMedBench (Qiu et al., 2024).
+Both evaluations underscore the superiority of StructTuning, surpassing other SOTA domain knowledge injection methods (Cheng et al., 2023; Zhang et al., 2024).
+Remarkably, we achieve a 50% improvement in knowledge injection compared to SOTA MMedLM2 in the medical domain, using only 0.3% of the training data requirement.
+Furthermore, StructTuning exhibits good scalability, achieving comparable performance with only 5% of the training data.
+These findings reveal our superiority in enhancing domain-specific AI assistants with more efficient data utilization.
+Figure 2: Framework for structure-aware knowledge injection.
+We extract the inherent knowledge structure in the training corpus, and associate training chunks to corresponding knowledge points.
+Models are continually pre-trained on data chunks in the condition of the knowledge structure, and fine-tuned with supervised QA samples to elicit their learned knowledge to solve knowledge-intensive (KI) and 2- or multi-hop questions in the real world.
 Our contribution is summarized as follows:
   • We proposed a novel two-stage training strategy, SCPT and SSFT, to inject domain knowledge into LLMs by preserving and utilizing the inherent structure of the training corpus.
-  • We developed a scalable data construction framework to generate structure-aware training samples from original corpora, so as to facilitate the SCPT and SSFT stages.
+  • We developed a scalable data construction framework to generate structure-aware training samples from original corpora to facilitate the SCPT and SSFT stages.
   • We conducted extensive investigations on our StructTuning strategy on various data and model settings, and comprehensively illustrate our superiority in knowledge injection.
-2 Related Work
-Domain Adaptation for Large Language Models.
-While pre-trained LLMs possess promising capabilities, their performance is often hampered by the scope and recency of their training data, which particularly affects smaller models in downstream applications (Zhao et al., 2023; Wang et al., 2023a).
-Continual Pre-Training (CPT) addresses this by perpetually updating a pre-trained model with domain-specific content (Sun et al., 2020; Xu et al., 2023b).
-To keep pace with the latest information, models can be fine-tuned with supervised instruction-response pairs (SFT), thus staying current with the advancing knowledge landscape (Mecklenburg et al., 2024; Qiu et al., 2024).
-Existing literature confirms that combining CPT and SFT is effective for LLMs to remain precise and up-to-date in dynamic fields like law (Cui et al., 2023; Nguyen, 2023) and medicine (Wang et al., 2023b; Qiu et al., 2024).
+2 Related Works
+Here we briefly discuss the closely related works. A detailed discussion can be found in Appendix C.
+Domain Adaptation for LLMs.
+To address the domain adaptation problem, a pre-trained model will be continually pre-trained (CPT) with domain-specific content (Sun et al., 2020; Xu et al., 2023b), and fine-tuned with supervised instruction-response pairs (SFT) to keep advancing interactive capabilities (Mecklenburg et al., 2024; Qiu et al., 2024).
+This paradigm is validated efective in dynamic fields like medicine (Wang et al., 2023b; Qiu et al., 2024) and coding (Roziere et al., 2023; Guo et al., 2024).
 Our study builds upon this CPT-SFT framework, innovating with SCPT-SSFT strategies to efficiently and effectively infuse domain knowledge with the inherent structure hierarchy.
 Structure-aware Knowledge Aggregation.
-A branch of researchers follows the conventional paradigm to extract entity-relation-entity triplets from texts to construct knowledge graphs (Pan et al., 2024), to enhance LLMs’s factual knowledge and logical reasoning by feature aggregation (Liu et al., 2020; Zhang et al., 2022), prompt engineering (Wen et al., 2023; Wang et al., 2023c), training data synthesis (Tang et al., 2024), etc.
-In these cases, each node corresponds to either a specific entity or an abstract concept, lacking the capability to present an informative and self-contained knowledge point.
-Some works have recently related a piece of descriptive text to a knowledge point, and constructed the knowledge structure for LLMs’ retrieval-augmented generation (Sarthi et al., 2024; Dong et al., 2024), where the top-to-down retrieval provides precise information-seeking paths along the knowledge structure.
-In this paper, we extend the structure-aware knowledge aggregation to LLMs’ training phase, injecting the whole domain knowledge structure into LLMs’ by linking the training samples into corresponding knowledge points and reasoning paths.
+In conventional paradigms, researchers extract entity-relation-entity triplets from texts to construct knowledge graphs (Pan et al., 2024), to enhance LLMs’s factual knowledge and logical reasoning(Zhang et al., 2022; Wen et al., 2023).
+Here, each node corresponds to either a specific entity or an abstract concept, lacking the capability to present an informative and self-contained knowledge point.
+This paper extends to structure-aware knowledge aggregation on existing training corpora, injecting the whole domain knowledge structure into LLMs’ by linking training samples to corresponding knowledge points and reasoning paths.
 Data Augmentation and Synthesis.
-Traditional methods aim to artificially expand the training dataset size (Xu et al., 2023a; Mukherjee et al., 2023) or generate entirely new samples that could help models learn better or adapt to specific tasks (Tang et al., 2024).
-Yet, they often overlook the structured nature of domain knowledge, and the aimlessly generated samples may also lack diversity (Ovadia et al., 2023; Mecklenburg et al., 2024), leading to potentially suboptimal training outcomes when applied for domain adaptations (Mecklenburg et al., 2024; Tang et al., 2024).
+Traditional methods aim to artificially expand the training dataset size (Xu et al., 2023a; Mukherjee et al., 2023) or generate entirely new samples to adapt LLMs to specific tasks (Tang et al., 2024).
+Yet, they often overlook the structured nature of domain knowledge, while the aimlessly generated samples may also lack diversity (Ovadia et al., 2023; Mecklenburg et al., 2024) and cannot cover the domain knowledge points (Mecklenburg et al., 2024; Tang et al., 2024).
 By contrast, our SSFT design is an innovative departure to address the challenge of retaining and utilizing the structured knowledge inherent in domain-specific content.
 3 Methodology
+Figure 3: Left: extracted knowledge structure. Right: template to bridge mindmap structure and textual contents.
 Fig.
 2 depicts our StructTuning methodology to inject domain knowledge into pre-trained LLMs using the inherent knowledge structure.
-With curated domain corpora (typically a few textbooks), we first develop an automatic approach to extract the knowledge structure, and associate text chunks to corresponding knowledge paths and points (Sec.
+With curated domain corpora (typically a few textbooks), we first extract the knowledge structure, and associate text chunks to corresponding knowledge paths and points (Sec.
 3.1).
@@
 3.3) techniques.
-Figure 2: Framework for structure-aware knowledge injection.
-We extract the inherent knowledge structure in the training corpus, and associate training chunks to corresponding knowledge points.
-Models are continually pre-trained on data chunks in the condition of the knowledge structure, and fine-tuned with supervised QA samples to elicit their learned knowledge to solve real-world questions (including knowledge-intensive (KI) QA, 2- or multi-hop QA, etc.).
-3.1 Automatic Extraction of Knowledge Structure
+3.1 Knowledge Structure Extraction
 For web-crawled corpus, previous data pre-processing focuses on quality assessment for individual documents (Bi et al., 2024), while the meta-info of knowledge structures (e.g., the table content for a textbook) is usually neglected or filtered out, and all we have are those sequentially arranged text segments (e.g., page-by-page-chunked content).
@@
 After that, we prompt the advanced Llama3-70B (Dubey et al., 2024) model to summarize the title for each chunk, where the textual content with the abstractive title jointly contributes to a “knowledge point”.
-Then, we automatically aggregate knowledge points and extract the inherent structure hierarchy by leveraging advanced language models.
+Then, we aggregate knowledge points and extract the inherent structure hierarchy by leveraging advanced language models.
 Inspired by Liu et al.
-(2024a), we take the title list to instruct a specifically developed 7B model to identify the inherent knowledge structure within the text chunks.
-A prompt example is displayed in Fig.
-3, and the detailed implementation is presented in Sec.
-B.1.
+(2024a), we take the title list to instruct a specifically developed 7B model to identify the inherent knowledge structure (as exemplified in Fig.
+3) within the text chunks, and Fig.
+A1 showcases how to deal with non-textbook data.
+The whole process does not involve human annotation, which reduces the cost and makes our method scalable for larger domain training corpora.
 In particular, Sec.
 B.1 and Sec.
-B.2 verify that our specialized 7B model can identify sufficiently precise knowledge structure for effective and efficient domain adaptation, as more powerful LLMs like LLaMA3-70B (Dubey et al., 2024) and GPT-3.5 (Brown et al., 2020) cannot bring significant enhancement while largely increase the inference costs.
-Fig.
-3 presents an example of the extracted structure, where we use the tree-like mindmap structure (Wen et al., 2023) to present the knowledge taxonomy from a textbook.
-And Fig.
-A1 showcases how to deal with non-textbook data.
-The whole process does not involve human annotation, which reduces the cost and makes our method scalable for larger domain training corpora.
-Figure 3: Left: prompt template to extract hierarchical knowledge structures from the given content. Right: example of the extracted knowledge structure presented by a mindmap format.
-After automatically extracting the domain knowledge structure and associating the original training chunks to related knowledge points, we delve into injecting domain knowledge through structure-aware continual pre-training (SCPT) and structure-aware supervised fine-tuning (SSFT).
+B.4 verify that our specialized 7B model can identify sufficiently precise knowledge structure for effective and efficient domain adaptation, as more powerful LLMs like LLaMA3-70B (Dubey et al., 2024) and GPT-3.5 (Brown et al., 2020) cannot bring significant enhancement while largely increase the inference costs.
 3.2 Structure-aware Continual Pre-Training
-In conventional knowledge injection methods, training corpora are randomly concatenated and chunked into text segments without distinguishing the original content, leading to the fact that models can only absorb domain knowledge that is emergent in the data diversity (Ovadia et al., 2023; Mecklenburg et al., 2024; Qiu et al., 2024).
-In this section, we present another solution to inject knowledge from limited pieces of textbooks by leveraging the highly abstractive and exhaustive domain knowledge structures for continual pre-training.
-Figure 4: Example of prompt templates to bridge mindmap structure and textual contents.
-We first transform the knowledge structure into natural languages using the same mindmap template in Fig.
-3, and prepend it to each training chunk, forcing LLMs to memorize the textual content (knowledge points) in the condition of the associated knowledge path in the structure hierarchy.
-We collected 20 diversified templates from GPT-4 (Achiam et al., 2023) to bridge mindmap structures and training chunks, one of which is displayed in Fig.
-4.
+Figure 4: QA samples synthesized for SSFT. We instruct Llama3-70B to generate (a) knowledge-intensive and (b) multi-hop questions and derive the diagnosis answers with explicit reasoning.
+In conventional knowledge injection, training corpora are randomly concatenated and chunked into text segments without distinguishing the original content, making models only able to absorb domain knowledge emerging in data diversity (Ovadia et al., 2023; Mecklenburg et al., 2024; Qiu et al., 2024).
+In this section, we present another solution to inject knowledge from limited text corpora by leveraging the highly abstractive and exhaustive domain knowledge structures for continual pre-training.
+We first transform the knowledge structure into natural languages using the same mindmap template Wen et al.
+(2023) in Fig.
+3 (left), and prepend it to each training chunk, forcing LLMs to memorize the textual content (knowledge points) in the condition of the associated knowledge path in the structure hierarchy.
+20 diversified templates (see Fig.
+A5) are collected from GPT-4 (Achiam et al., 2023) to bridge mindmap structures and training chunks, one of which is displayed in Fig.
+3 (right).
 The prepended mindmap, as well as the template, does not produce auto-regressive loss.
@@
 Formally, we turn the original language modeling in vanilla CPT to conditioned modeling (Keskar et al., 2019) in our SCPT stage:
-$p({\bm{x}}^{k})=\prod_{i=1}^{n}{p(x^{k}_{i}|x^{k}_{<i})}\quad\Longrightarrow\quad p({\bm{x}}^{k}|{\bm{s}}^{k})=\prod_{i=1}^{n}{p(x^{k}_{i}|x^{k}_{<i},{\bm{s}}^{k})}$ (1)
-where $p({\bm{x}}^{k})$ models the probability distribution for the $k$-th chunk ${\bm{x}}^{k}=(x^{k}_{1},\cdots,x^{k}_{n})$ via the chain rule of probability (Bengio et al., 2000) on each token $x^{k}_{i}$, and ${\bm{s}}^{k}$ denotes the associated knowledge mindmap paths.
-As illustrated in Fig.
-2 (b), after traversing the $m$ knowledge points in extracted structures, models are asked to recall the whole knowledge hierarchy, i.e., to model the composed probability distribution:
+$p({\bm{x}}^{k})=\prod_{i=1}^{n}{p(x^{k}_{i}|x^{k}_{<i})}\quad\Longrightarrow$ (1)
+$p({\bm{x}}^{k}|{\bm{s}}^{k})=\prod_{i=1}^{n}{p(x^{k}_{i}|x^{k}_{<i},{\bm{s}}^{k})}$
+where $p({\bm{x}}^{k})$ models the probability distribution for the $k$-th chunk ${\bm{x}}^{k}=(x^{k}_{1},\cdots,x^{k}_{n})$ via the chain rule of probability (Bengio et al., 2000) on each token $x^{k}_{i}$, and ${\bm{s}}^{k}$ denotes the associated knowledge mindmap.
+Sec.
+B.5 extensively investigates the effectiveness of our SCPT strategy.
+As Fig. 2 (b) shows, after traversing the $m$ knowledge points in extracted structures, models are asked to recall the whole knowledge hierarchy, i.e., to model the composed probability distribution:
 $p(\bar{{\bm{s}}})=\prod_{k=1}^{m}{p({\bm{s}}^{k})}$ (2)
@@
 These two steps iteratively alternate throughout training epochs.
-Next, we will introduce how to teach LLMs to explicitly utilize their domain knowledge, which is learned in the SCPT stage, to solve practical problems by doing exercises with our SSFT technique.
 3.3 Structure-aware Supervised Fine-Tuning
-In traditional knowledge injection paradigms, supervised fine-tuning serves as a tool to align the (continually) pre-trained models to interactive chatbots, e.g., through massive question-answering exercises (Cui et al., 2023; Qiu et al., 2024).
-However, most QA data augmentation or generation strategies aim at enlarging the quantity and enhancing the diversity of training samples (Xu et al., 2023a; Mukherjee et al., 2023; Liu et al., 2024b), which neglects the nature of the highly structured domain knowledge.
-Therefore, our structure-aware supervised fine-tuning (SSFT) technique focuses on eliciting models’ structured knowledge learned during the SCPT stage, adapting LLMs to interactive and reliable domain experts.
+Traditional supervised fine-tuning aims to align the (continually) pre-trained models to interactive ChatBots through question-answering exercises (Cui et al., 2023; Qiu et al., 2024).
+Previous studies focus on enlarging the quantity and enhancing the diversity of training syntheses (Xu et al., 2023a; Mukherjee et al., 2023; Liu et al., 2024b) but neglect the highly structured domain knowledge.
+In contrast, our structure-aware supervised fine-tuning (SSFT) technique aims to elicit models’ structured knowledge learned during SCPT, adapting LLMs to interactive and reliable domain experts.
 Fig.
-2 (c) illustrates the idea of synthesizing SSFT samples guided by domain knowledge structures.
-Specifically, we use the random walk algorithm to create knowledge paths with $1$ to $l$ branches in the original mindmap.
-For paths linking to a single knowledge point, we use the corresponding text content to prompt Llama3-70B (Dubey et al., 2024) to generate knowledge-intensive question-answering pairs.
+2 (c) illustrates SSFT samples synthesis guided by domain knowledge structures.
+First, we use the random walk algorithm to create knowledge paths with $1$ to $l$ branches in the original mindmap (the illustration of knowledge paths and branches is displayed in Fig.
+A2).
+For paths linking to a single knowledge point, we use the corresponding text content to prompt Llama3-70B (Dubey et al., 2024) to generate knowledge-intensive QA pairs.
 For paths with two or more branches, we prompt Llama3-70B with the knowledge path and textual contents to synthesize 2- or multi-hop QA samples, which require specific reasoning along the knowledge structure to derive from questions to answers.
@@
 4 presents several examples.
-Figure 5: QA samples synthesized for SSFT. We instruct Llama3-70B to generate (a) knowledge-intensive and (b) multi-hop questions and derive the diagnosis answers with explicit reasoning.
 For every synthesized QA sample (${\bm{z}}$), we will prepend the relevant mindmap hierarchy to the answer, and add a CoT prompt in the question to construct another type of QA data (${\bm{z}}^{\prime}$) for SFT alignment.
@@
 (2024).
-During testing, we can either use the vanilla question as input to efficiently gather models’ answers, or take the CoT prompt to probe to what extent LLMs can memorize and leverage the injected knowledge to answer the questions.
-Integrating with SCPT and SSFT, our StructTuning approach translates into remarkable efficacy and efficiency in domain knowledge injection, as comprehensively evaluated in the following sections.
+During testing, we use the vanilla question as input to efficiently gather models’ answers to calculate accuracy, and take the CoT prompt to probe to what extent LLMs can memorize and leverage the injected knowledge to answer the questions.
+Integrating with SCPT and SSFT, our StructTuning approach translates into remarkable efficacy and efficiency in domain knowledge injection, as comprehensively evaluated in subsequent sections.
 4 Experiments
-We design a comprehensive evaluation of our StructTuning through several experiments on two benchmarks.
-First, we investigate the free-form question-answering task on the LongBench (Bai et al., 2023) dataset, in order to verify the memorization and understanding of injected knowledge (the answer can be directly found in training corpora).
-Then, we delve into the multi-choice question-answering task on MMedBench (Qiu et al., 2024), to explore how LLMs apply the injected knowledge in basic medicine to determine the real-world diagnosis for patients with logical reasoning.
-4.1 Preliminary Investigation on Free-form Question-Answering
-Table 1: Recall evaluation of Open-Book QA (OBQA) and Closed-Book QA (CBQA) tasks on the LongBench (Bai et al., 2023) dataset.
-The best results are marked in bold, and the secondary results are marked with underlines.
-The backbone model is Llama2-7B (Touvron et al., 2023b).
+Table 1: Recall of Closed-Book QA (CBQA) on LongBench (Bai et al., 2023b). The best and secondary results are marked in bold and underlines, respectively. The base model is Llama2-7B.
 Task Adaptation SingleDoc-QA MultiDoc-QA Average
 Qasper MFQA MFQAzh HpQA 2Wiki Musiq Duzh
-OBQA - 39.7 44.3 17.5 30.1 35.5 11.9 9.6 26.9
 CBQA CPT+SFT 20.7 35.3 20.6 29.9 32.1 18.9 12.0 24.2
@@
 SCPT+SSFT 30.5 44.6 24.3 40.8 42.0 21.8 16.8 31.5
+We design extensive evaluations of our StructTuning through several experiments on two benchmarks.
+First, we investigate the free-form question-answering task on the LongBench (Bai et al., 2023b) dataset, so as to verify the memorization and understanding of injected knowledge (the answer can be directly found in training corpora).
+Then, we delve into the multi-choice question-answering task on MMedBench (Qiu et al., 2024), to explore how LLMs apply the injected knowledge in basic medicine to determine the real-world diagnosis for patients with logical reasoning.
+4.1 Preliminary Free-form QA Investigation
 Datasets and Tasks.
-LongBench (Bai et al., 2023) is a multi-task benchmark tailored for open-book reading comprehension evaluation, where LLMs generate answers to given questions based on one or several input passages.
-To focus on knowledge injection, we turn the open-book evaluation into a closed-book QA task, where LLMs are trained on contextual passages and queried with questions only.
-We choose 7 subsets with 1,350 test examples from LongBench for single- and multi-document question-answering evaluation, and the remaining synthetic or code-orientated tasks are eliminated.
-More details are described in Sec.
+Seven subsets with 1,350 test examples from LongBench (Bai et al., 2023b) are utilized to evaluate closed-book question-answering evaluation, where the answers can be directly found in corresponding passages.
+The 14K reading-comprehension passages are used for knowledge injection via CPT/SCPT, and another 2,700 QA samples are generated for SFT/SSFT.
+Details are described in Sec.
 A.1.
 Evaluation Metrics.
-To quantify the knowledge memorization degree, we mainly report the recall (Zhu & Li, 2023b) for models’ outputs against ground-truth answers for the free-form QA tasks.
+Here we first report the recall (Zhu and Li, 2023b) for models’ outputs against ground-truth answers to quantify the knowledge memorization degree.
 In Sec.
-B.3 we also evaluate our method with the F1-score measure for a thorough comparison.
-Investigated Models. For LongBench, we mainly investigate the knowledge injection to Llama2-7B (Touvron et al., 2023b) to compare the open- and closed-book QA performance.
-Implementation Details.
-For closed-book QA, the Llama2-7B model is continually pre-trained on 10,476 passages for 3 epochs using a batch size of 128.
-The initial learning rate is 2e-5, decayed to 0 at the end of training with a cosine scheduler.
-Then, we query Llama3-70B to generate 2,700 SSFT QA examples and remove those with over 0.5 F1-Score similarity to test samples to prevent data leakage.
-We train the knowledge-injected models for 1 epoch to avoid overfitting.
+B.9 we also evaluate F1-score for a thorough comparison.
+Investigated Models. We mainly investigate Llama2-7B (Touvron et al., 2023b) to compare the knowledge injection performance.
+Implementation Details. We train all models on LongBench passages for 3 epochs using a batch size of 128, and train for 1 epoch on synthetic SFT data to avoid overfitting. The learning rate is 2e-5.
+Main Results.
+We first try to inject passage content into LLMs with a conventional CPT+SFT paradigm for the baseline, and use CoT instructions during testing to elicit models’ memorized knowledge in their responses.
+However, Tab.
+1 indicates such an injection approach is ineffective, as the knowledge recall is only 24.2%.
+On the other hand, our SCPT strategy achieved a higher knowledge recall of 26.7%.
+It implies the model has successfully associated the relevant passages with their entire knowledge structure for the given question, especially on multi-doc QAs that require more complex information retrieval and reasoning on multiple passages.
+Furthermore, our SSFT technique continues to improve the knowledge recall to 31.5%, which indicates the vanilla SFT strategy can only regularize LLMs’ response styles, while our SSFT could teach LLMs to utilize their knowledge (injected in the SCPT stage) to answer corresponding questions.
+Fig.
+5 showcases a representative example.
 Figure 5: An example for structure-aware responses.
-Main Results.
-First, we use a CoT instruction of “Answer the question and explain why” to set a baseline for Open-Book QA (OBQA).
-As shown in Tab.
-1, it presents a moderate knowledge comprehending capability in SingleDoc QA tasks (e.g., over 40% recall on Qasper and MFQA), and relatively poor knowledge coordination in MultiDoc QA subsets (e.g., around 10% recall on Musiq and Duzh) due to the attention drift on longer inputs.
-Then, we try to inject passage content into LLMs with a conventional CPT+SFT paradigm to benchmark the Closed-Book QA (CBQA) baseline.
-Note that we use two types of question templates (refer to Sec.
-3.3) during training, and use the CoT version during testing to elicit models’ memorized knowledge in their responses.
-However, Tab.
-1 indicates such a knowledge injection approach is ineffective, as the knowledge recall is even lower than the OBQA baseline (i.e., 24.2% v.s.
-26.9%).
-On the contrary, our SCPT strategy has already achieved a considerable knowledge recall of 26.7%, which is higher than the CPT-SFT paradigm of 24.2% and approaching the OBQA baseline of 26.9%.
-It implies our structure-aware continually pre-trained model has successfully associated the relevant passages with their entire knowledge structure for the given question, which provides the knowledge path to seek targeted information to derive the answer.
-Furthermore, our SSFT technique explicitly teaches the model to recall the learned knowledge and answer the questions, which further improves the knowledge recall to 31.5% and even surpasses the Open-Book QA setting.
-The results indicate the vanilla SFT strategy can only regularize LLMs’ response styles, while our SSFT could teach LLMs to utilize their knowledge (injected in the SCPT stage) to answer corresponding questions, as exemplified in Fig.
-6.
-Particularly, our method receives significant enhancements on MultiDoc QA subsets, which implies even though the total text content may exceed the models’ attention window and influence the OpenBook-QA, we can still chunk the content into several pieces and successfully inject the knowledge points into LLMs meanwhile preserving the whole knowledge structure.
-Figure 1: MindMap Recall
-F1-Score BERTScore
-0.61 0.87
-In addition, we also use lexical ROUGE-L (Lin, 2004) and semantic BERTScore (Zhang et al., 2020) to quantify the memorization of injected knowledge structures, by comparing the mindmap in models’ responses (as Fig.
-6 displays) with ground-truth answers.
-The results in Fig.
-1 indicate a relatively good memorization of the injected knowledge mindmap, emphasizing the efficacy of our SCPT strategy.
-4.2 In-depth Evaluation for Multi-choice QA Application
+In Sec. B.8, we also validate relatively good memorization of injected knowledge structures, further emphasizing our SCPT’s efficacy.
+Table 2: Multiple-choice evaluation on MMedBench (Qiu et al., 2024).
+We report separate accuracies across six languages, with “Average” denoting the mean score.
+“#Token” denotes required data for knowledge injection.
+Model English Chinese Japanese French Russian Spanish Average #Token
+ChatDoctor (Yunxiang et al., 2023) 43.52 43.26 25.63 18.81 62.50 43.44 39.53 -
+PMC-LLaMA Wu et al. (2024) 47.53 42.44 24.12 20.74 62.11 43.29 40.04 -
+MedAlpaca Han et al. (2023) 46.74 44.80 29.64 21.06 59.38 45.00 41.11 -
+Llama2-7B Touvron et al. (2023b) 43.36 50.29 25.13 20.90 66.80 47.10 42.26 -
+InternLM2-7B Zheng et al. (2024) 57.27 77.55 47.74 41.00 68.36 59.59 58.59 -
+Llama3-8B (AI, 2024) 63.86 78.23 48.24 50.80 71.48 64.15 62.79 +0.00 -
+Llama3+MMed Qiu et al. (2024) 66.06 79.25 61.81 55.63 75.39 68.38 67.75 +4.96 25.5B
+Llama3+StructTuning (Ours) 66.77 77.44 53.27 51.61 74.61 68.49 65.36 +2.57 76M
+Llama3+StructTuning (Ours) 65.36 79.04 56.28 55.47 80.47 69.80 67.74 +4.95 1.2B
+4.2 In-depth Multi-choice QA Evaluation
 Datasets and Tasks.
-MMedBench (Qiu et al., 2024) is a multilingual medical multi-choice QA benchmark, with 45,048 QA pairs for adapting LLMs to medical experts and 8,518 for testing.
-We collect 76M textbook corpora from MedTextBooks (Jin et al., 2020) and MMedC (Qiu et al., 2024) for medical knowledge injection and use the training/test split from MMedBench for SFT/evaluation.
-We also curate another two sizes of training sets (with 30M and 132M tokens) to validate our method’s scalability.
+We take several corpus sizes from MMedC (Qiu et al., 2024) for CPT/SCPT, with 45K QA data from MMedBench (Qiu et al., 2024)’s training set for SFT/SSFT.
+Models are evaluated on six multi-choice subsets from MMedBench, where LLMs should make real-world patient diagnoses with adequate reasoning on medical knowledge.
 Detailed setup is in Sec.
@@
 Evaluation Metrics.
-For multi-choice QA in MMedBench (Qiu et al., 2024), we follow the default setting to calculate the accuracy on six language subsets, as well as the averaged scores.
+We follow the default setting to calculate the accuracy on six language subsets and the averaged scores.
 Metrics are computed by lexical exact-matching on models’ responses, rather than maximum token probabilities.
 Investigated Models.
-For MMedBench, we extend the investigated LLMs across model scales and architectures including Llama2-7B/13B (Touvron et al., 2023b), InternLM2-7B (Zheng et al., 2024), and the recent Llama3-8B (Dubey et al., 2024).
-We also compare our knowledge-injected models with other popular domain-specified LLMs, such as MedAlpaca (Han et al., 2023), ChatDoctor (Yunxiang et al., 2023), PMC-LLaMA (Wu et al., 2024), and MMedLM (Qiu et al., 2024) models.
+We extend the investigation across model architectures and scales, including Llama2-7B/13B (Touvron et al., 2023b), InternLM2-7B (Zheng et al., 2024), and Llama3-8B (Dubey et al., 2024).
+Other popular medical LLMs (Han et al., 2023; Wu et al., 2024; Qiu et al., 2024) are also included for a thorough comparison.
 Implementation Details.
-We first train LLMs for 3 epochs on medical textbooks with a batch size of 128.
-Then, we ask Llama3-70B to create structure-aware explanations for existing 45K QA samples in MMedBench’s training split and 33K extra entries by traversing extracted knowledge structures.
-Syntheses with overlapped options in the test set are removed.
-In the SFT phase, the learning rate is set as 1e-6 to avoid overfitting on such an amount of SFT samples, as suggested by Qiu et al.
-(2024).
+Following Qiu et al.
+(2024), models are first trained for 3 epochs on medical corpora with a learning rate of 2e-5, and then fine-tuned for 1 epoch to avoid overfitting.
+The detailed setup is displayed in Sec.
+A.2.
+Figure 6: Knowledge injection’s scalability.
 Main Results.
-In Tab.
-2, we present the overall performance across a series of LLMs on the testing split of MMedBench.
-The results demonstrate the promising enhancement achieved by our StructTuning technique, which translates into consistent improvements on the advanced InternLM2-7B (Zheng et al., 2024) and Llama3-8B (Dubey et al., 2024) models, and largely outperforms the previous domain-specific LLMs like PMC-LLaMA (Wu et al., 2024) and MedAlpaca (Han et al., 2023).
-Notably, our structure-aware knowledge injection approach, using merely 76M tokens curated from medical textbooks, achieves over 50% performance (4.46% v.s.
-8.71%, 2.57% v.s.
-4.96%) against the state-of-the-art MMedLM (Qiu et al., 2024) method, which is trained on a huge MMedC (Qiu et al., 2024) corpora of 25.5B tokens.
-Our approach shows much more efficiency in transforming pre-trained LLMs into domain experts by structured-aware knowledge injection, e.g., we only use 0.3% corpus to achieve comparable or even slightly better performance on English and Spanish subsets on top of Llama3-8B.
-However, our method is currently unable to achieve 100% of knowledge injection efficacy against MMedLM, which may comprise two major reasons: (1) some knowledge in the training corpus has already been learned by the foundation model (e.g., a slight decrease in the Chinese subset), and (2) knowledge in the curated 76M training data cannot cover all testing scenarios (e.g., a considerable gap on the Japanese subset due to the absence of Japanese corpus, see Tab.
-A2).
-Therefore, we make a step to evaluate our approach’s scalability on various corpora sizes in Fig.
-8.
-Table 2: Multiple-choice accuracy evaluation on MMedBench (Qiu et al., 2024).
-We report each model’s accuracy across six languages separately, with “Average” denoting the mean score over six languages.
-We also compare the data requirement (“#Token”) for medical knowledge injection.
-Model English Chinese Japanese French Russian Spanish Average #Token
-ChatDoctor 43.52 43.26 25.63 18.81 62.50 43.44 39.53 -
-PMC-LLaMA 47.53 42.44 24.12 20.74 62.11 43.29 40.04 -
-MedAlpaca 46.74 44.80 29.64 21.06 59.38 45.00 41.11 -
-Llama2-7B 43.36 50.29 25.13 20.90 66.80 47.10 42.26 -
-InternLM-7B 44.07 64.62 37.19 24.92 58.20 44.97 45.67 -
-InternLM2-7B 57.27 77.55 47.74 41.00 68.36 59.59 58.59 +0.00 -
-+MMed 61.74 80.01 61.81 52.09 80.47 67.65 67.30 +8.71 25.5B
-+Ours 60.80 79.19 50.75 45.34 75.39 66.85 63.05 +4.46 76M
-Llama3-8B 63.86 78.23 48.24 50.80 71.48 64.15 62.79 +0.00 -
-+MMed 66.06 79.25 61.81 55.63 75.39 68.38 67.75 +4.96 25.5B
-+Ours 66.77 77.44 53.27 51.61 74.61 68.49 65.36 +2.57 76M
-Figure 8: Comparison of knowledge injection approaches.
+The results in Tab.
+2 demonstrate the promising enhancement achieved by our StructTuning technique largely outperforming the previous domain-specific LLMs like PMC-LLaMA (Wu et al., 2024) and MedAlpaca (Han et al., 2023).
+Notably, our structure-aware knowledge injection approach, using merely 76M tokens (0.3%) curated from medical textbooks, achieves over 50% performance (2.57% v.s.
+4.96%) against the state-of-the-art MMedLM (Qiu et al., 2024) method, which is trained on the entire MMedC (Qiu et al., 2024) corpora of 25.5B tokens.
+After scaling up the training tokens to 1.2B (around 5%), our method makes nearly 100% improvements to the average accuracy, significantly reducing the training cost of traditional knowledge injection approaches.
 Approach’s Scalability.
-In Tab.
-2, we use roughly 0.3% of the 25.5B tokens in MMedC Qiu et al.
-(2024) to evaluate knowledge injection methods.
-Here, another two training corpora sizes are curated: 30M and 132M, which take around 0.1% and 0.5% of 25.5B tokens respectively.
-We compare the vanilla CPT-SFT paradigm and our SCPT-SSFT strategy across those data settings The backbone LLM is InternLM2-7B (Zheng et al., 2024), and the training settings follow the main experiments.
+We further curate a series of training corpora sizes to investigate our method’s scalability in-depth: 30M, 76M, 132M, 250M, and 1.2B, which respectively take around 0.1%, 0.3%, 0.5%, 1%, and 5% of 25.5B tokens.
+The vanilla CPT-SFT paradigm and our SCPT-SSFT strategy are thoroughly compared across those data settings.
 According to Fig.
-8, our method consistently surpasses the vanilla paradigm by a large margin, emphasizing the efficacy and efficiency of domain knowledge injection.
-In particular, we can fit two performance-ratio scaling curves from the data points in Fig.
+6, our method consistently surpasses the vanilla paradigm by a large margin, emphasizing the efficacy and efficiency of domain knowledge injection.
+In particular, we fit two performance-ratio scaling curves from the data points in Fig.
 6 as:
-$p_{v}\approx-0.04(\log r)^{2}+13.3\log r+100.0;\quad p_{s}\approx-1.11(\log r)^{2}+7.63\log r+133.0$ (3)
-where $p_{v}$ and $p_{s}$ denote the relative performance enhancement (%) for vanilla and structure-aware knowledge injection, and $r$ refers to the corpus ratio.
-The scaling law indicates we can achieve comparable performance (100%) with only 5% of the total training corpus, significantly reducing the costs for LLMs’ domain adaptation.
-On the other hand, it also indicates our method may lead to 133% enhancement with a further 100% comprehensive data utilization.
-However, we have not empirically verified those predictions due to time and resource constraints, which we view as a temporal limitation of our current work and leave it to future investigations.
-Approach’s Generalization.
-In addition to the InternLM2 and Llama3 models in previous experiments, we also investigate the generalization ability of our SCPT+SSFT paradigm on Llama2 (Touvron et al., 2023b) model series.
-As shown in Tab.
-3, our method leads to consistently significant improvements on 7B (+8.78%) and 13B (+6.17%) backbone models.
-The results further demonstrate the generalizability and scalability of our StructTuning strategy across model architectures and sizes.
-Table 3: Structure-aware knowledge injection to Llama2 (Touvron et al., 2023b) model series.
-Model English Chinese Japanese French Russian Spanish Average
-Llama2-7B 43.36 50.29 25.13 20.90 66.80 47.10 42.26
-+Ours 49.41 65.15 36.68 35.21 69.14 50.62 51.04
-Llama2-13B 51.37 57.97 32.66 25.08 69.92 52.99 48.33
-+Ours 53.02 68.30 37.78 41.71 70.70 55.51 54.50
-Ablation Studies.
-To further demonstrate the efficacy of each component of our StructTuning method, we conduct a comprehensive ablation study with the English split of the MMedBench dataset.
-Specifically, we take Llama2-7B as the backbone model, select the English textbooks (Jin et al., 2020) (with 26M tokens) for vanilla and our structure-aware continual pre-training, and use different QA samples for supervised fine-tuning.
-In Tab.
-4, “SFT” refers to vanilla SFT with 10K English QA samples provided in MMedBench’s (Qiu et al., 2024) training split, “SSFT” indicates structure-aware SFT on 10K training data, where the questions are the same as “SFT” while the answers are enhanced with knowledge structure explanation by Llama3-70B, as described in Sec.
-3.3.
-“SSFT*” includes another 8K structure-aware QA syntheses by Llama3-70B, consisting of 18K entries for training.
-The training hyper-parameters follow the main experiment.
+$p_{v}\approx-0.04(\log r)^{2}+13.3\log r+100.0$ (3)
+$p_{s}\approx-1.11(\log r)^{2}+7.63\log r+133.0$
+where $p_{v}$ and $p_{s}$ denote the relative performance enhancement (%) for vanilla and structure-aware knowledge injection, and $r$ is the corpus ratio.
+In fact, we use the scaling law derived from 0.1%, 0.3%, 0.5%, and 1% points to predict that achieving 100% performance would require 5% of the data corpus, which has been confirmed in Tab.
+2.
+On the other hand, it also indicates our method may lead to 133% enhancement with a further 100% comprehensive data utilization, further validating the effectiveness and scalability of our method.
+Table 3: Generalization to various model architectures and sizes. Here we use 76M training corpus.
+Model Size Adaptation Accuracy
+InternLM2 7B CPT+SFT 58.59
+SCPT+SSFT 63.05
+Llama2 7B CPT+SFT 42.26
+SCPT+SSFT 51.04
+Llama2 13B CPT+SFT 48.33
+SCPT+SSFT 54.50
 Table 4: Ablation studies with Llama2-7B on the English subset of MMedBench.
@@
 SFT 44.54 32.81 26.63 15.27 53.91 42.30 35.91
-CPT SFT 46.27 32.57 26.13 17.36 50.00 40.63 35.49
-SCPT SFT 46.50 32.14 20.10 18.17 53.91 39.97 35.13
-SCPT SSFT 49.96 32.63 22.11 17.52 51.17 41.28 35.78
-SCPT SSFT* 49.10 33.92 18.33 27.14 57.42 43.73 38.27
+CPT + SFT 46.27 32.57 26.13 17.36 50.00 40.63 35.49
+SCPT + SFT 46.50 32.14 20.10 18.17 53.91 39.97 35.13
+SCPT + SSFT 49.96 32.63 22.11 17.52 51.17 41.28 35.78
+SCPT + SSFT* 49.10 33.92 18.33 27.14 57.42 43.73 38.27
 RAG 38.12 29.22 22.61 23.34 53.91 36.47 33.95
+AdaptLLM (Cheng et al., 2023) 46.79 33.80 20.60 14.15 53.12 42.34 35.03
+RAFT (Zhang et al., 2024) 43.60 32.34 21.11 14.95 50.39 42.16 34.09
+Approach’s Generalization.
 In Tab.
-4, vanilla CPT+SFT injection paradigm does bring an improvement of 1.73% in the English split, while our SCPT technique with vanilla SFT strategy presents a slightly higher accuracy of 46.50%.
-However, when combining SSFT with SCPT, our method immediately leads to a significant boost in the English split (49.96% v.s.
-44.54%), demonstrating the efficacy and necessity of eliciting the learned structured knowledge to solve practical problems.
-Moreover, the supplementation of 8K extra QA syntheses (“SSFT*”) surprisingly enhances the model performance on the other five subsets, resulting in a valuable gain on the average accuracy to 38.27%.
-This result demonstrates the knowledge transferability across different languages (Lai et al., 2023; Qin et al., 2024).
-After training with SSFT, LLMs can actively utilize the knowledge injected in one language to solve the problem in another language, evidencing our superiority against the traditional SFT technique.
-In Sec.
-B.4, we follow Liu et al.
-(2024b) to randomly generate another 8K SFT pairs for further comparison, where our structure-aware SFT syntheses are verified to better enhance LLMs’ knowledge application.
+3, we also validated the performance on Llama2 (Touvron et al., 2023b) and InternLM2 (Zheng et al., 2024) model series by using 76M tokens for knowledge injection.
+Our method leads to consistently significant improvements on InternLM2-7B (+4.46%), Llama2-7B (+8.78%) and Llama2-13B (+6.17%) backbone models, further demonstrating the generalizability and scalability of our StructTuning across model architectures and sizes.
+Detailed results are presented in Tab.
+A6.
+Ablation Studies.
+We perform a comprehensive ablation study on MMedBench’s English subset in Tab.
+4.
+As suggested by Qiu et al.
+(2024), we use the English textbooks (Jin et al., 2020) (26M tokens) to compare vanilla and structure-aware CPT, paired with corresponding SFT strategies.
+In particular, “SFT” uses vanilla SFT with 10K QA samples from MMedBench’s training split, while “SSFT” applies structure-aware SFT on the same questions, enhancing answers with Llama3-70B-generated knowledge explanations (Sec.
+3.3).
+“SSFT*” further includes 8K additional structure-aware QA pairs, totaling 18K training entries.
+Training hyperparameters align with the main experiment.
+In Tab.
+4, the CPT+SFT paradigm improves accuracy by 1.73%, while SCPT with vanilla SFT achieves a higher 46.50%.
+Combining SCPT with SSFT boosts performance significantly (49.96% v.s.
+44.54%), highlighting the importance of structured knowledge elicitation.
+Adding 8K extra QA pairs (“SSFT*”) further improves performance across five subsets, demonstrating a surprising cross-lingual knowledge transfer (Lai et al., 2023; Qin et al., 2024).
+After SSFT, LLMs effectively use knowledge injected in one language to solve problems in others, surpassing traditional SFT.
+Additional comparisons in Sec.
+B.6 confirm that structure-aware syntheses can enhance knowledge application better than random syntheses.
 In addition, we observe that the commonly used RAG (Lewis et al., 2020) strategy does not bring significant advantages to the MMedBench evaluation.
@@
 Knowledge injection by (S)CPT and (S)SFT shows more advantages in this situation.
+In-depth investigations can be found in Sec.
+B.7.
+Comparision with Other Methods.
+We also compare two advanced knowledge injection methods in Tab.
+4 to further demonstrate our StructTuning’s efficacy: (1) AdaptLLM (Cheng et al., 2023): domain knowledge injection by appending reading comprehension QAs to each CPT chunk, and (2) RAFT (Zhang et al., 2024): improving LLM’s robustness to domain-specific retrieval-augmented generation using noisy retrieval-augmented SFT samples.
+According to the experimental results, AdaptLLM (Cheng et al., 2023) brings negligible improvement in the final performance (e.g., 46.79% v.s.
+46.27% on the English subset), indicating such a chunk-level reading comprehension augmentation during CPT cannot help LLMs capture the entire structured domain knowledge.
+Concurrently, RAFT (Zhang et al., 2024) causes even worse performance, since the retrieval process introduces too many unrelated chunks and hurts LLM’s QA judgments, especially when there exists a significant gap between user query and knowledge chunks in the medical diagnosis scenario.
 5 Conclusion
@@
 Through a novel SCPT-SSFT paradigm, we have set a new precedent for adapting LLMs to specialized domains, and the promising and scalable results underscore the viability and potential of our method.
-We hope to inspire further research in efficient and effective domain adaptation, moving a step closer to models that can truly emulate human intelligence.
+We hope to inspire further research in efficient and effective domain adaptation in the LLM community, moving a step closer to models that can truly emulate human intelligence.
 Limitation.
-A noteworthy limitation of our work is the under-explored full potential of our StructTuning strategy for knowledge injection.
-Although we have estimated a scaling law in Eq.
-3, the data requirement to achieve 100% of the desired effectiveness remains unverified.
-Besides, our method introduces negligible data preprocessing costs to extract the domain knowledge structure.
-However, Sec.
-B.2 indicates our method can still reduce the overall adaptation cost to 10% since only a small part of training corpus is required.
-We will delve into the investigations in our future work.
+Our two-stage strategy introduces added computational complexity, where taxonomy extraction and data reorganization are required in the SCPT phase, and extra QA syntheses are optionally applied in the SSFT stage.
+In Appendix B, we provide further discussion with extensive empirical experiments.
+Despite the additional computational overhead introduced, our method achieves greater overall benefits and can reduce the reliance on large-scale LLMs (e.g., 70B models).
+We will delve into the investigations in future work.
 Appendix A Implementation Details
 A.1 Detailed Setup on LongBench
-To focus on the investigation of knowledge injection, we choose 7 subsets from LongBench (Bai et al., 2023) across single- and multi-document QA tasks in English and Chinese, and the remaining synthetic or code-orientated tasks are eliminated:
+Dataset Composition.
+To focus on the investigation of knowledge injection, we choose 7 subsets from LongBench (Bai et al., 2023b) across single- and multi-document QA tasks in English and Chinese, and the remaining synthetic or code-orientated tasks are eliminated:
   • Single-Doc QA.
-For single-document QA, we take three subsets from LongBench: (1) Qasper (Dasigi et al., 2021), featured by question-answering over NLP technical papers and annotated by NLP practitioners; (2) MultiFieldQA (Bai et al., 2023), manually curated from multiple data sources and annotated by Ph.D.
+For single-document QA, we take three subsets from LongBench: (1) Qasper (Dasigi et al., 2021), featured by question-answering over NLP technical papers and annotated by NLP practitioners; (2) MultiFieldQA (Bai et al., 2023b), manually curated from multiple data sources and annotated by Ph.D.
 students; and (3) MultiFieldQA-zh, the Chinese split also provided by Bai et al.
-(2023), covering multiple Chinese scenarios.
-MultiFieldQA contains 150 Context-Question-Answer triplets to test, and the others adopted subsets include 200 pieces of test samples respectively.
+(2023b), covering multiple Chinese scenarios.
+MultiFieldQA contains 150 Context-Question-Answer triplets to test, and the others subsets include 200 pieces of test samples respectively.
   • Multi-Doc QA.
@@
 We take four multi-hop QA datasets: (1) HotpotQA (Yang et al., 2018), containing 2-hop questions written by native speakers given two related paragraphs; (2) 2WikiMultihopQA (Ho et al., 2020), involving up to 5-hop questions synthesized through manually designed templates on Wikipedia passages; (3) MuSiQue (Trivedi et al., 2022), carefully composed with up to 4-hop reasoning on an increased number of supporting and distracting context evidence; and (4) Dureader (He et al., 2017), developed based on Baidu Search and Baidu Zhidao and filtered by Bai et al.
-(2023) to reduce the data noise.
+(2023b) to reduce the data noise.
 Each subset has 200 test samples.
-For data entries in Single-Doc QA, we extract the knowledge structure for each single passage; in Multi-Doc QA, we identify the knowledge structure across multiple passages for each test sample.
+In Single-Doc QA, we extract knowledge structures for each single passage; in Multi-Doc QA, we identify the knowledge structure across multiple passages for each test sample.
 There are ultimate 1350 question-answer-passage(s)-(knowledge)structure quadruples to evaluate knowledge injection approaches on LongBench.
+SFT Data-Synthesis.
+We query Llama3-70B to generate 2,700 QA examples and remove those with over 0.5 F1-Score similarity to test samples to prevent data leakage.
+During inference, when the model can generate correct answers (corresponding to specific knowledge points) that haven’t been seen during the SFT stage, we can ensure the knowledge is injected at the CPT stage and SFT only enhances the instruction-following capability.
+In practice, merely 13 out of 2700 (around 0.5%) synthetic data have over 0.5 F1-Score and are thus filtered out from the SFT data.
+Tab.
+A1 statistics the semantic similarity (measured by BERTScore (Zhang et al., 2020)) between generated and GT questions and answers, and the results emphasize there is no knowledge leakage in the generated SFT data (they share poor semantic similarity across questions, answers, and QAs).
+Table A1: Similarity statistics on synthetic SFT data and LongBench’s test samples.
+Target Question Answer Question-Answer
+BERTScore 0.277 0.106 0.093
+Table A2: Sample statistics on MMedBench’s QA data.
+Split English Chinese Japanese French Russian Spanish Total
+Train 10,178 27,400 1,590 2,171 1,052 2,657 45,048
+Test 1,273 3,426 199 622 256 2,742 8,518
+Table A3: Sample statistics on different training data settings.
+Stage Ratio English Chinese Japanese French Russian Spanish Total
+CPT 0.1% 6.9M 8.8M - 4.1M 4.9M 5.4M 30.1M
+CPT 0.3% 26.1M 21.5M - 8.1M 10.3M 10.1M 76.1M
+CPT 0.5% 35.9M 27.2M 4.5M 14.0M 24.9M 26.1M 132.6M
+CPT 1.0% 44.1M 39.8M 34.2M 45.1M 47.9M 38.4M 249.5M
+CPT 5.0% 169.4M 227.8M 119.8M 232.7M 235.5M 226.1M 1.2B
+SFT - 18.8K 39.1K 1.6K 5.3K 5.9K 7.5K 78.2K
 A.2 Detailed Setup on MMedBench
@@
 Statistics on six languages are displayed in Tab.
-A1.
+A2.
 Notably, the benchmark includes scenarios where questions may have multiple correct answers (i.e., in Japanese and French subsets), introducing additional complexity for the model evaluation process.
-Table A1: Sample statistics on MMedBench.
-Split English Chinese Japanese French Russian Spanish Total
-Train 10,178 27,400 1,590 2,171 1,052 2,657 45,048
-Test 1,273 3,426 199 622 256 2,742 8,518
 Data for Continual Pre-Training.
-To investigate high-quality domain knowledge injection for LLMs, we collect 18 English textbooks and 33 Chinese textbooks from the National Medical Board Examination in the USA and Mainland China, respectively (Jin et al., 2020).
-All collected textbooks are originally in PDF format and Jin et al.
-(2020) converted them into digital text via OCR and performed some clean-up pre-processing strategies to reduce the data noise.
-The English and Chinese textbooks count for around 26.1M and 21.5M tokens by Llama2 tokenizer (Touvron et al., 2023b).
-Then, we randomly sample an extra 28M textbook corpora from MMedC (Qiu et al., 2024) for the other languages (except the unavailable Japanese textbooks), and the statistics are displayed in Tab.
-A2.
-Table A2: Sample statistics on training data. $\dagger$” and $\ddagger$” refer to different sizes.
-Stage English Chinese Japanese French Russian Spanish Total
-CPT 26.1M 21.5M - 8.1M 10.3M 10.1M 76.1M
-[…]
-[…]
-SFT 18.8K 39.1K 1.6K 5.3K 5.9K 7.5K 78.2K
-[…]
+To investigate high-quality domain knowledge injection for LLMs and the scalability of injection methods, we curate a series of training corpus sizes from the 25.5B MMedC (Qiu et al., 2024) dataset, including 0.1%, 0.3%, 0.5%, 1%, and 5%, which respectively takes 30M, 76M, 132M, 250M, and 1.2B tokens.
+We sorted the document-level text content (including textbooks and other corpus from websites and wikipedia) by length in descending order and progressively included more samples to expand the training dataset at larger scales, where Tab.
+A3 provides detailed statistics.
+In particular, as MMedC does not provide English textbooks in its released data (due to copyright issues), we collect 18 English textbooks (Jin et al., 2020) as an alternative for English medical knowledge injection, since they have a common OCR source with MMedC Qiu et al.
+(2024).
+These English textbooks take around 21.5M tokens.
+Knowledge Structure Extraction.
+For textbooks, we split the data to chunks (knowledge points) within 3072 tokens, and use our specifically developed 7B-size LLM to extract the structured medical knowledge system.
+For non-textbook data (for instance, MMedC does not provide Japanese textbooks), a clustering-based technique (Sarthi et al., 2024) is adopted to recursively build knowledge structures from fragmented text segments.
+Fig.
+A1 presents an example to illustrate the two kinds of knowledge structure extraction processes, Tab.
+A4 displays the overall statistics on extracted knowledge structures for the final 1.2B tokens.
 Data for Supervised Fine-Tuning.
@@
 The final quantity statistics are presented in Tab.
-A2.
+A3.
+Note that the 70B-size model is not necessary to synthesize QAs and explanations, since we have provided relevant text sources (retrieved from the training corpus) to supplement medical knowledge.
+In this way, LLMs only need to perform in-context comprehension, rather than generate new QAs based on their own knowledge, which can also be achieved by smaller models like Qwen2.5-7B Yang et al.
+(2024).
+Table A4: Knowledge structure on 1.2B training corpus.
+Lang. Book Chapter Section KnowledgePoints
+6 2933 14,411 23,239 180,793
+A.3 Terminology Explanation
 Knowledge Structures.
@@
 Here the terms of “Section”, “Chapter”, and “book” are just examples to help illustrate the knowledge structure.
-A.3 Resource Requirement
+Figure A2: Definition and example of knowledge paths and branches.
+Knowledge Paths and Branches. Fig. A2 shows an example of how we define the knowledge paths and branches of the extracted knowledge structure for SSFT data synthesis.
+  • A path means a knowledge path from the domain summary (e.g., Biochemistry) to specific knowledge points (e.g., Lipid Metabolism and Cholesterol Transport): “Biochemistry – Overview of lipoprotein metabolism, hormone synthesis – Lipoprotein Metabolism – Lipid Metabolism and Cholesterol Transport”.
+  • A branch means the knowledge branch of the tree structure.
+If a question is related to two knowledge points (e.g., Lipid Metabolism and Cholesterol Transport and Pancreatic zymogen activation) at different branches of the knowledge tree, the knowledge path contains two branches, which becomes the right-bottom part of Fig.
+A2.
+A.4 Resource Requirement
 We use 8 NVIDIA A100-80G GPUs to train all the models, and leverage 1-2 NVIDIA A100-80G GPUs for inference.
 Appendix B Additional Experiments
-B.1 Ablation on Knowledge Structure Extraction
+B.1 Knowledge Structure Extraction
 Extracting domain knowledge structure is a prerequisite for subsequent knowledge injection (including both SCPT and SSFT) for language models.
@@
 3.1, we propose a bottom-up strategy to re-chunk the texts from domain textbooks, summarize a title for each chunk, and send the title list to a specialized 7B model to derive the knowledge structure.
-The prompt template is displayed in Appendix C.
+The prompt template is displayed in Appendix D.
 In fact, we argue that due to the language diversity, a perfectly recovered table of contents of textbooks is unnecessary for domain knowledge injection.
@@
 In Tab.
-A3, we individually adopt few-shot GPT-3.5-Turbo (Brown et al., 2020) and LLaMA3-70B (Dubey et al., 2024) models to extract medical knowledge structure from 18 English textbooks (Jin et al., 2020) (with 26.1M tokens) for subsequent knowledge injection (the backbone LLM is LLaMA2-7B (Touvron et al., 2023b)).
+A5, we individually adopt few-shot GPT-3.5-Turbo (Brown et al., 2020) and LLaMA3-70B (Dubey et al., 2024) models to extract medical knowledge structure from 18 English textbooks (Jin et al., 2020) (with 26.1M tokens) for subsequent knowledge injection (the backbone LLM is LLaMA2-7B (Touvron et al., 2023b)).
 Although they present 3.6%-3.7% enhancement on MMedBench (Qiu et al., 2024)’s English test set (denoted as “improvement”), leveraging GPT-3.5-Turbo and LLaMA3-70B is either expensive or time-consuming.
-GPT-3.5-Turbo costs around 15 dollars to process 26M tokens, while LLaMA3-70B takes around 1.5 hours on 2 A100-80G GPUs, of which both limit the scalability of data pre-processing for structure-aware knowledge injection.
+GPT-3.5-Turbo costs around 15 dollars to process 26M tokens, while LLaMA3-70B takes around 1.5 hours on 2 A100-80G GPUs, of which both limit scaling data pre-processing for structure-aware knowledge injection.
 Table A5: Comparison of models to extract knowledge structures on 26M English corpus.
-Model Improvement Time Consumption Extra Cost
+Model Improvement Time Cost Extra Fee
 GPT-3.5-Turbo +3.58 - 15\$
@@
 Ours-7B +3.69 0.2h -
+Table A6: Structure-aware knowledge injection to Llama2 (Touvron et al., 2023b) model series.
+Model English Chinese Japanese French Russian Spanish Average
+InternLM2-7B 57.27 77.55 47.74 41.00 68.36 59.59 58.59
++Ours 60.80 79.19 50.75 45.34 75.39 66.85 63.05
+Llama2-7B 43.36 50.29 25.13 20.90 66.80 47.10 42.26
++Ours 49.41 65.15 36.68 35.21 69.14 50.62 51.04
+Llama2-13B 51.37 57.97 32.66 25.08 69.92 52.99 48.33
++Ours 53.02 68.30 37.78 41.71 70.70 55.51 54.50
 Inspired by Liu et al.
-(2024a), we distilled the knowledge structure extraction capability from giant LLMs to a LLaMA2-7B model via supervised fine-tuning.
+(2024a), we distilled the knowledge structure extraction capability from giant LLMs to a LLaMA2-7B model via SFT.
 In particular, we instruct LLaMA3-70B to generate 22K training examples (pairs of raw knowledge points and extracted knowledge structures) from Wikipedia, and train a LLaMA2-7B model at a batch size of 128 and a learning rate of 2e-5 for 1 epoch.
 After utilizing the specialized 7B model to identify the knowledge structure in medical textbooks, as shown in Tab.
-A3, the results translate to comparable performance on structure-aware knowledge injection.
-Meanwhile, the inference cost significantly decreases to 0.3 hours, which is more scalable to handle a larger amount of domain corpus.
-B.2 Comparison on Training Costs
+A5, the results translate to comparable performance on structure-aware knowledge injection.
+Meanwhile, the inference cost significantly decreases to 0.3 hours, which is more scalable to handle a larger domain corpus.
+Table A7: Model evaluation on general benchmarks.
+Injection MMLU C-Eval TruthfulQA WinoGrande ARC_c
+Before 0.46 0.35 0.49 0.52 0.51
+After 0.43 0.35 0.43 0.51 0.56
+B.2 Evaluation of Approach’s Generalization
+In addition to the Llama3 models in previous experiments, we also investigate the generalization ability of our SCPT+SSFT paradigm on Llama2 (Touvron et al., 2023b) and InternLM2 (Zheng et al., 2024) model series.
+As shown in Tab.
+A6, our method leads to consistently significant improvements on InternLM2-7B (+4.46%), Llama2-7B (+8.78%) and Llama2-13B (+6.17%) backbone models.
+The results further demonstrate the generalizability and scalability of our StructTuning strategy across model architectures and sizes.
+B.3 Investigation on Model Overfitting
+Here we provide a supplementary evaluation on common benchmarks in Tab.
+A7, which indicates the current methodology of structure-aware knowledge injection does not significantly hurt LLM’s general capabilities, and even brings slight enhancement to the ARC_c benchmark (maybe the knowledge structure enhances the reasoning ability).
+As previous research states, the issue of overfitting can be further mitigated by incorporating common QA examples, and we leave this in our future work.
+B.4 Comparison on Training Costs
+Table A8: Training costs for knowledge injection.
+Paradigm Corpus Improve. Preprocess Total
+CPT+SFT 100% 100% - $>$30d
+SCPT+SSFT 0.3% 50% 0.6h 4.5h
+SCPT+SSFT 5% 100% 9.7h 3d
+Table A9: Ablation studies of SCPT on MMedBench subsets. The base model is Llama2-7B.
+Adaptation English Chinese Japanese French Russian Spanish Average
+CPT+SFT 46.27 32.57 26.13 17.36 50.00 40.63 35.49
+Ours-FixTmpl1 48.27 32.86 20.61 23.70 56.17 42.56 37.36
+Ours-FixTmpl2 48.10 32.99 21.23 23.97 55.97 43.10 37.56
+Ours-RemoveL1 47.90 32.90 20.11 24.63 57.10 43.43 37.68
+Ours-RemoveL2 48.05 33.63 21.62 24.11 57.49 43.13 38.00
+Ours-RemoveL3 48.47 33.14 20.15 23.33 56.86 43.65 37.60
+Ours-NTPLoss 48.99 33.15 20.57 25.31 56.78 42.94 37.96
+Ours-Full 49.10 33.92 18.33 27.14 57.42 43.73 38.27
 In Tab.
-A4, we quantify the total training cost on 8 A100-80G GPUs.
+A8, we quantify the total training cost on 8 A100-80G GPUs.
 According to Qiu et al.
@@
 As suggested in Fig.
-8, when 5% training data is leveraged for knowledge injection to achieve 100% improvement, the overall cost is limited to 3 days, much less than the CPT+SFT approach with more than a month.
+6, when 5% training data is leveraged for knowledge injection to achieve 100% improvement, the overall cost is limited to 3 days, much less than the CPT+SFT approach with more than a month.
 Those analyses further demonstrate the efficacy and efficiency of our structure-aware knowledge injection framework.
-Table A4: Comparison of training costs for knowledge injection.
-Paradigm Corpus Improvement Pre-process Total
-CPT+SFT 100% 100% - $>$30d
-SCPT+SSFT 0.3% 50% 0.6h 4.5h
-SCPT+SSFT 5% 100% 9.7h 3d
-B.3 F1-Score Evaluation on LongBench
-In Sec.
-4.1, we mainly follow Zhu & Li (2023b) to investigate the memorization and understanding of injected knowledge by calculating the knowledge recall in models’ responses.
-Here we report the F1-score measure over the Open-Book QA (OBQA) and Closed-Book QA (CBQA) settings for a thorough comparison.
-Note that here we use the vanilla question prompt to obtain concise answers, instead of the CoT prompt used in Sec.
-4.1 to elicit models’ memorized knowledge.
-The evaluated models are the same as Sec.
-4.1.
-In Tab.
-A5, we report the Open-Book QA (OBQA) baseline for Llama2-7B with passages as inputs, which shows the best performance on MultiFieldQA (MFQA) (Bai et al., 2023) and 2WikiMultihopQA (2Wiki) (Ho et al., 2020) subsets.
-Then, we establish the Closed-Book QA (CBQA) baseline by traditional CPT+SFT to inject passage contents into model parameters, and supplement the experiment of our SCPT+SSFT technique for comparison.
-According to the results shown in Tab.
-A5, CPT+SFT slightly improves the QA performance on several subsets (such as MultiFieldQA-zh (MFQAzh) (Bai et al., 2023)), while the overall F1-Score measure is still inferior to the OBQA performance.
-In contrast, our SCPT+SSFT approach successfully boosts the closed-book QA performance to 20.1% on average, even surpassing the open-book QA baseline of 18.7%.
-The results are consistent with Sec. 4.1, which jointly demonstrate the effectiveness of structure-aware knowledge injection for large language models.
-Table A5: F1 Score evaluation of Open-Book QA (OBQA) and Closed-Book QA (CBQA) tasks on the LongBench (Bai et al., 2023) dataset.
-The best results are marked in bold, and the secondary results are marked with underlines.
-The backbone model is Llama2-7B (Touvron et al., 2023b).
-Task Adaptation SingleDoc-QA MultiDoc-QA Average
-Qasper MFQA MFQAzh HpQA 2Wiki Musiq Duzh
-OBQA - 19.2 36.8 11.9 25.4 32.8 9.4 5.2 18.7
-CBQA CPT+SFT 16.8 23.1 13.2 21.3 19.1 10.4 13.4 16.8
-SCPT+SFT 15.2 21.5 15.2 14.9 19.8 5.6 14.1 15.2
-SCPT+SSFT 19.7 23.5 19.5 26.4 24.1 12.2 15.4 20.1
-B.4 Ablation on SFT Data Synthesis
+B.5 Ablation on Structured Knowledge Injection
+During the Structure-aware Continual Pre-Training (SCPT) stage, we proposed to learn specific text chunks (knowledge points) in the condition of the mindmap inputs (knowledge structures), in order to relate the knowledge points to corresponding structure nodes.
+In this section, we conduct a series of ablation studies to investigate the design efficacy.
+The vanilla CPT+SFT paradigm is adopted as the comparison baseline, where the Llama2-7B model is trained with CPT and SFT data on the English subset of MMedBench, while tested on all subsets across six languages.
+The hyper-parameter settings follow the main experiment in our manuscript.
+The empirical results are presented in Tab.
+A9.
+First, we investigate the choice of formatting template to convert the knowledge structure to a mindmap condition.
+In particular, we try to fix the template to convert all knowledge mindmaps for SCPT, and randomly select two templates to repeat the experiment.
+According to Tab.
+A9, fixed SCPT templates lead to inferior performance against randomly choosing the template from the diversified 20 template pool.
+This is consistent with Zhu and Li (2023a)’s observation, that text rewriting can provide better knowledge augmentation for large language models.
+Then, we explore the impact of the extracted knowledge structure itself.
+In MMedBench, a 3-layer knowledge structure (follow the chapter-section-subsection hierarchy) is constructed for each textbook, and we respectively remove the 1st (chapter), 2nd (section), and 3rd (subsection) layer of the hierarchy during knowledge injection.
+As Tab.
+A9 shows, removing the top layer (chapter) leads to the worst performance of 47.90%, because the remaining knowledge points cannot effectively relate to each other without the organization of the top layer.
+On the other hand, removing the bottom layer (subsection) performs slightly better on the English subset (because of the controlled structure-information lost), but hinders the cross-language knowledge utilization on the remaining subsets (e.g., 37.60% on average across six languages).
+Finally, we revisit the modeling choice of the mindmap-conditioning learning.
+Specifically, we try to turn the conditional modeling $p({\bm{x}}^{k}|{\bm{s}}^{k})$ back to complete next-token prediction $p({\bm{x}}^{k},{\bm{s}}^{k})$ (the next-token prediction loss is computed on mindmap condition as well).
+According to Tab.
+A9, the performance is slightly inferior to our full version of SCPT strategy (e.g., 37.96% v.s.
+38.27% on Average).
+Therefore, we reserve conditional modeling for our SCPT stage.
+B.6 Ablation on SFT Data Synthesis
+Table A10: Comparison of SFT data synthesis strategies on MMedBench. The backbone LLM is the same Llama2-7B model after SCPT on English textbooks.
+SFT synthesis English Chinese Japanese French Russian Spanish Average
+- 46.50 32.14 20.10 18.17 53.91 39.97 35.13
+SFT* 47.13 32.49 16.58 16.72 51.95 42.16 34.51
+SSFT* 49.10 33.92 18.33 27.14 57.42 43.73 38.27
 In Sec.
@@
 Tab.
-A6 indicates that “SFT*” brings slight enhancement to the English test subset, but the average accuracy drops to 34.51% instead.
+A10 indicates that “SFT*” brings slight enhancement to the English test subset, but the average accuracy drops to 34.51% instead.
 The results further demonstrate our method’s efficacy in the application of the injected, structured domain knowledge.
-Table A6: Comparison of SFT data synthesis strategies on MMedBench. The backbone LLM is the same Llama2-7B model after SCPT on English textbooks.
-SFT synthesis English Chinese Japanese French Russian Spanish Average
-- 46.50 32.14 20.10 18.17 53.91 39.97 35.13
-SFT* 47.13 32.49 16.58 16.72 51.95 42.16 34.51
-SSFT* 49.10 33.92 18.33 27.14 57.42 43.73 38.27
-Appendix C Prompt Template for Knowledge Structure Extraction
+B.7 In-depth Comparison on Retrieval-Augmented Generation
+Table A11: Ablation on the hyper-parameter settings for the RAG baseline.
+ChunkSize 256 512 1024
+RetrieveNum 10 5 3 5 3 2 3 2 1
+Accuracy 35.08 37.67 38.04 36.42 38.12 37.99 35.00 36.89 38.07
+Table A12: Attempts to integrate hybrid-search and reranker models.
+Hybrid Reranker Accuracy
+$\times$ $\times$ 38.12
+$\surd$ $\times$ 37.97
+$\times$ $\surd$ 37.52
+$\surd$ $\surd$ 37.75
+In Sec.
+4.2, we briefly compare RAG adaptation and injection-based approaches in the MMedBench dataset, and this section provides more implementation details and further investigations on the popular retrieval-augmented generation approach.
+Experimental Settings.
+On the implementation of the RAG baseline, we utilize the BAAI/bge-m3 (Chen et al., 2024) embedding model for dense retrieval, due to its state-of-the-art and multi-lingual semantic retrieval ability.
+For the experiments in Tab.
+4, we take the same 26M English CPT data as the knowledge base, re-chunk the data corpus for every 512 tokens, and retrieve top-3 related chunks as context inputs for LLM’s generation process.
+The retrieval process is implemented using the LlamaIndex [^2] https://www.llamaindex.ai/ framework.
+Additional Experiments.
+We also conduct a variety of experiments to evaluate the hyper-parameters for the RAG baseline.
+As shown in Tab.
+A11, changing the chunk size and retrieved chunk number cannot bring any significant benefits.
+The core reason lies in the gap between user query and retrieved chunks.
+In particular, user queries contain many descriptive and quantitative sentences and numbers (such as the example in Fig.
+A3, “They enrolled 800 patients in the study, half of which have breast cancer”.), and may even talk about an entirely new thing that has not been recorded in the knowledge base.
+Furthermore, we also try to use the hybrid (dense+sparse) search strategy and larger rerank model (BAAI/bge-reranker-v2-m3 (Chen et al., 2024)) to enhance the retrieval quality.
+However according to the results in Tab.
+A12, the semantic gap between user queries and retrieved chunks still exists.
+Introducing the hybrid search and rerank model even gets worse performance (e.g., the keyword age may be considered a key factor for hybrid search, but it cannot help to derive the answer of test sensitivity).
+Conclusion.
+RAG may assist in some knowledge-intensive tasks for information-seeking, but will encounter problems when there exists a significant semantic gap between user query and retrieved documents.
+MMedBench is a typical scenario, where LLMs are asked to derive medical diagnoses with proper reasoning according to the descriptions of patients or medical examinations.
+In this case, the retrieval process introduces too many unrelated chunks and hurts LLM’s QA judgments.
 Fig.
-A2 displays the prompt template to query our specialized 7B model to extract knowledge structure on given knowledge points, which introduces the task definition, detailed instruction, and output formats to illustrate the process.
+A3 provides an example where the retrieved chunks are actually unrelated to the complicated user query (the user asks about the analysis of a given research study, but the retrieved documents contain several keywords, e.g., age, while having nothing to do with the blood test study.)
+Figure A3: An example of retrieved document/chunk based on a given query.
+B.8 Knowledge Memorization on LongBench
+Here, we also use lexical ROUGE-L (Lin, 2004) and semantic BERTScore (Zhang et al., 2020) to quantify the memorization of injected knowledge structures, by comparing the mindmap in models’ responses (as Fig.
+5 displays) with ground-truth answers.
+The results in Tab.
+A13 indicate a relatively good memorization of the injected knowledge mindmap, emphasizing the efficacy of our SCPT strategy.
+Table A13: MindMap Recall
+F1-Score BERTScore
+0.61 0.87
+B.9 F1-Score Evaluation on LongBench
+Table A14: F1 Score evaluation of Closed-Book QA (CBQA) tasks on the LongBench (Bai et al., 2023b) dataset.
+The best results are marked in bold, and the secondary results are marked with underlines.
+The backbone model is Llama2-7B (Touvron et al., 2023b).
+Task Adaptation SingleDoc-QA MultiDoc-QA Average
+Qasper MFQA MFQAzh HpQA 2Wiki Musiq Duzh
+CBQA CPT+SFT 16.8 23.1 13.2 21.3 19.1 10.4 13.4 16.8
+SCPT+SFT 15.2 21.5 15.2 14.9 19.8 5.6 14.1 15.2
+SCPT+SSFT 19.7 23.5 19.5 26.4 24.1 12.2 15.4 20.1
+In Sec.
+4.1, we mainly follow Zhu and Li (2023b) to investigate the memorization and understanding of injected knowledge by calculating the knowledge recall in models’ responses.
+Here we report the F1-score measure over the Closed-Book QA (CBQA) settings for a thorough comparison.
+Note that here we use the vanilla question prompt to obtain concise answers, instead of the CoT prompt used in Sec.
+4.1 to elicit models’ memorized knowledge.
+The evaluated models are the same as Sec.
+4.1.
+In Tab.
+A14, we report the Closed-Book QA (CBQA) baseline by traditional CPT+SFT to inject passage contents into model parameters, and supplement the experiment of our SCPT+SSFT technique for comparison.
+According to the results shown in Tab.
+A14, our SCPT+SSFT approach successfully boosts the closed-book QA performance to 20.1% on average.
+The results are consistent with Sec.
+4.1, which jointly demonstrate the effectiveness of structure-aware knowledge injection for large language models.
+Appendix C Detailed Related Work
+Domain Adaptation for LLMs.
+While pre-trained LLMs possess promising capabilities, their performance is often hampered by the scope and recency of their training data, which particularly affects smaller models in downstream applications (Zhao et al., 2023; Wang et al., 2023a).
+Continual Pre-Training (CPT) addresses this by perpetually updating a pre-trained model with domain-specific content (Sun et al., 2020; Xu et al., 2023b).
+, with parameter-efficient tuning methods devised to curtail training costs (Hu et al., 2021; Liu et al., 2024c).
+To keep pace with the latest information, models can be fine-tuned with supervised instruction-response pairs (SFT), thus staying current with the advancing knowledge landscape (Mecklenburg et al., 2024; Qiu et al., 2024).
+Existing literature confirms that combining CPT and SFT is effective for LLMs to remain precise and up-to-date in dynamic fields like medicine (Wang et al., 2023b; Qiu et al., 2024) and coding (Roziere et al., 2023; Guo et al., 2024).
+Our study builds upon this CPT-SFT framework, innovating with SCPT-SSFT strategies to efficiently and effectively infuse domain knowledge with the inherent structure hierarchy.
+Conditional Language Modeling.
+The idea of continual pre-training language models on domain corpus in the condition of the knowledge structure is mainly inspired by CTRL (Keskar et al., 2019).
+Keskar et al.
+(2019) demonstrates the effectiveness of steering text generation through control codes (one or two words) that signify the desired genre, style, or task.
+In the era of LLM, system prompt plays a similar role in controlling models’ responses to adapt to different needs and functionalities, such as role-playing, language style transfer, task setting, and behavior setting (Brown et al., 2020; Wang et al., 2023d; Bai et al., 2023a).
+Our SCPT approach extends the control codes or system prompts to domain-specific knowledge structures, so as to guide the learning process and tailor the model’s output more closely to specialized fields.
+Structure-aware Knowledge Aggregation.
+Knowledge structure has been widely explored in the recent LLM community.
+In conventional paradigms, researchers extract entity-relation-entity triplets from texts to construct knowledge graphs (Pan et al., 2024), to enhance LLMs’s factual knowledge and logical reasoning by feature aggregation (Liu et al., 2020; Zhang et al., 2022), prompt engineering (Wen et al., 2023; Wang et al., 2023c), information searching (Logan IV et al., 2019), training data synthesis (Tang et al., 2024), etc.
+In these cases, each node corresponds to either a specific entity or an abstract concept, lacking the capability to present an informative and self-contained knowledge point.
+Some works have recently related a piece of descriptive text to a knowledge point, and constructed the knowledge structure for LLMs’ retrieval-augmented generation (Sarthi et al., 2024; Dong et al., 2024), where the top-to-down retrieval provides precise information-seeking paths along the knowledge structure.
+In this paper, we extend the structure-aware knowledge aggregation to LLMs’ training phase, injecting the whole domain knowledge structure into LLMs’ by linking training samples to corresponding knowledge points and reasoning paths.
+Data Augmentation and Synthesis.
+Due to the lack of high-quality datasets, data augmentation has emerged as a promising solution to mimic real-world patterns (Liu et al., 2024b).
+Traditional methods aim to artificially expand the training dataset size (Xu et al., 2023a; Mukherjee et al., 2023) or generate entirely new samples that could help models learn better or adapt to specific tasks (Tang et al., 2024).
+Yet, they often overlook the structured nature of domain knowledge, and the aimlessly generated samples may also lack diversity (Ovadia et al., 2023; Mecklenburg et al., 2024), leading to potentially suboptimal training outcomes when applied for domain adaptations (Mecklenburg et al., 2024; Tang et al., 2024).
+By contrast, our SSFT design is an innovative departure to address the challenge of retaining and utilizing the structured knowledge inherent in domain-specific content.
+Appendix D Prompt Template for Knowledge Structure Extraction
+Fig.
+A4 displays the prompt template to query our specialized 7B model to extract knowledge structure on given knowledge points, which introduces the task definition, detailed instruction, and output formats to illustrate the process.
 Figure A4: Prompt template for knowledge structure identification.
+Figure A5: Full template pool for mindmap conversion with 20 diversified templates.